### v1

# Variational Flow Matching for Graph Generation

Abstract
We present a formulation of flow matching as variational inference, which we refer to as variational flow matching (VFM). Based on this formulation we develop CatFlow, a flow matching method for categorical data. CatFlow is easy to implement, computationally efficient, and achieves strong results on graph generation tasks. In VFM, the objective is to approximate the posterior probability path, which is a distribution over possible end points of a trajectory. We show that VFM admits both the CatFlow objective and the original flow matching objective as special cases. We also relate VFM to score-based models, in which the dynamics are stochastic rather than deterministic, and derive a bound on the model likelihood based on a reweighted VFM objective. We evaluate CatFlow on one abstract graph generation task and two molecular generation tasks. In all cases, CatFlow exceeds or matches performance of the current state-of-the-art models.
**footnotetext: These authors contributed equally to this work.
1 Introduction
In recent years, the field of generative modeling has seen notable advancements. In image generation [35, 38], the development and refinement of diffusion-based approaches — specifically those using denoising score matching [51] — have proven effective for generation at scale [14, 46]. However, while training can be done effectively, the constrained space of sampling probability paths in a diffusion requires tailored techniques to work [44, 56]. This is in contrast to more flexible approaches such as continuous normalizing flows (CNFs) [8], that are able to learn a more general set of probability paths than diffusion models [45], at the expense of being expensive to train as they require one to solve an ODE during each training step (see e.g. [5, 39, 13]).
Recently, Lipman and collaborators [24] proposed flow matching (FM), an efficient and simulation-free approach to training CNFs. Concretely, they use a per-sample interpolation to derive a simpler objective for learning the marginal vector field that generates a desired probability path in a CNF. This formulation provides equivalent gradients without explicit knowledge of the (generally intractable) marginal vector field. This work has been extended to different geometries [7, 20] and various applications [52, 9, 12, 21]. Similar work has been proposed concurrently in [27, 1].
This paper identifies a reformulation of flow matching that we refer to as variational flow matching (VFM). In flow matching, the vector field at any point can be understood as the expected continuation toward the data distribution. In VFM, we explicitly parameterize the learned vector field as an expectation relative to a variational distribution. The objective of VFM is then to minimize the Kullback-Leibler (KL) divergence between the posterior probability path, i.e. the distribution over possible end points (continuations) at a particular point in the space, and the variational approximation.
We show that VFM recovers the original flow matching objective when the variational approximation is Gaussian and the conditional vector field is linear in the data. Under the assumption of linearity, a solution to the VFM problem is also exact whenever the variational approximation matches the marginals of the posterior probability path, which means that we can employ a fully-factorized variational approximation without loss of generality.
While VFM provides a general formulation, our primary interest in this paper is its application to graph generation, where the data are categorical. This setting leads to a simple method that we refer to as CatFlow, in which the objective reduces to training a classifier over end points on a per-component basis. We apply CatFlow to a set of graph generation tasks, both for abstract graphs [30] and molecular generation [34, 18]. By all metrics, our results match or substantially exceed those obtained by existing methods.
2 Background
2.1 Transport Framework for Generative Modeling and CNFs
Common generative modeling approaches such as normalizing flows [36, 32] and diffusion models [14, 46] parameterize a transformation $\varphi$ from some initial tractable probability density $p_{0}$ – typically a standard Gaussian distribution – to the target data density $p_{1}$. In general, there is a trade-off between allowing $\varphi$ to be expressive enough to model the complex transformation while ensuring that the determinant term is still tractable. One such transformation is a continuous normalizing flow (CNF).
Any time-dependent [^1] Time dependence is denoted through the subscript $t$ throughout this paper. vector field $v_{t}:[0,1]\times\mathbb{R}^{D}\to\mathbb{R}^{D}$ gives rise to such a transformation – called a flow – as such a field induces a time-dependent diffeomorphism $\varphi_{t}:[0,1]\times\mathbb{R}^{D}\to\mathbb{R}^{D}$ defined by the following ordinary differential equation (ODE):
$\frac{d}{dt}\varphi_{t}(x)$ $=v_{t}(\varphi_{t}(x))\text{ with initial conditions }\varphi_{0}(x)=x.$ (1)
In CNFs, this vector field is learned using a neural network $v_{t}^{\theta}$. Through the change of variables formula $p_{t}(x)$ can be evaluated (see appendix D) and hence one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution. However, obtaining a gradient sample for the loss requires ones to solve the ODE induced during training, making this approach computationally expensive.
2.2 Flow Matching
In flow matching [24], the aim is to regress the underlying vector field of a CNF directly on the interval $t\in[0,1]$. Flow matching leverages the fact that even though we do not have access to the actual underlying vector field – which we denote as $u_{t}$ – and probability path $p_{t}$, one can construct a per-example formulation by defining conditional flows, i.e. the trajectories towards specific datapoints $x_{1}$. Concretely, FM sets:
$u_{t}(x)=\int u_{t}(x\mid x_{1})\frac{p_{t}(x\mid x_{1})p_{\text{data}}(x_{1})}{p_{t}(x)}\mathrm{d}x_{1},$ (2)
where $u_{t}(x\mid x_{1})$ is the conditional trajectory. The most common way to define $u_{t}(x\mid x_{1})$ is as the straight line continuation from $x$ to $x_{1}$, implying one can obtain samples $x\sim p_{t}(x\mid x_{1})$ simply by interpolating samples $x_{0}\sim p_{0}$ for some $p_{0}$ and $x_{1}\sim p_{1}$,
$u_{t}(x\mid x_{1}):=\frac{x_{1}-x}{1-t}\implies x=tx_{1}+(1-t)x_{0}\text{ is asample }x\sim p_{t}(x\mid x_{1}).$ (3)
Crucially, the flow matching objective
$\mathcal{L}_{\text{FM}}(\theta)=\mathbb{E}_{t\sim[0,1],x\sim p_{t}(x)}\left[\big{\|}v_{t}^{\theta}(x)-u_{t}(x)\big{\|}_{2}^{2}\right]$ (4)
is equivalent in expectation (up to a constant) to the conditional flow matching objective
$\mathcal{L}_{\text{CFM}}(\theta)=\mathbb{E}_{t\sim[0,1],x_{1}\sim p_{\text{data}}(x_{1}),x\sim p_{t}(x\mid x_{1})}\left[\big{\|}v_{t}^{\theta}(x)-u_{t}(x\mid x_{1})\big{\|}_{2}^{2}\right].$ (5)
An advantage of flow matching is that this conditional trajectory $u_{t}(x\mid x_{1})$ can be chosen to make the problem tractable. The authors show that diffusion models can be instantiated as flow matching with a specific conditional trajectory, but also show that assuming a simple, straight-line trajectory leads to more efficient training. Note that in contrast to likelihood-based training of CNFs, flow matching is simulation free, leading to a scalable approach to learning CNFs.
3 Variational Flow Matching for Graph Generation
We derive CatFlow through a novel, variational view on flow matching we call Variational Flow Matching (VFM). The VFM framework relies on two insights. First, we can define the marginal vector field and its approximation in terms of an expectation with respect to a distribution over end points of the transformation. This implies that we can map a flow matching problem onto a variational counterpart, opening up the usage of variational inference techniques. Second, under typical assumptions on the forward process, we can decompose the expected conditional vector field into components for individual variables which can be computed in terms of the marginals of the distribution over end points of the conditional trajectories. This implies that, without loss of generality, we can solve a VFM problem using a fully-factorized variational approximation, providing a tractable approximate vector field. For categorical data, the corresponding vector field can be computed efficiently via direct summation. This results in a closed-form objective to train CNFs for categorical data, which we refer to as CatFlow. Note that we develop the theory of VFM in section 3 and we relate VFM to flow matching and score-based diffusion in section 4.
3.1 Flow Matching using Variational Inference
In any flow matching problem, the vector field in eq. 2 can be expressed as an expectation
$u_{t}(x)$ $=\int u_{t}(x\mid x_{1})p_{t}(x_{1}\mid x)\mathrm{d}x_{1}=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right],$ (6)
where $p_{t}(x_{1}\mid x)$ is the posterior probability path, the distribution over possible end points $x_{1}$ of paths passing through $x$ at time $t$,
$p_{t}(x_{1}\mid x)$ $:=\frac{p_{t}(x,x_{1})}{p_{t}(x)},$ $p_{t}(x,x_{1})$ $:=p_{t}(x\mid x_{1})\>p_{\text{data}}(x_{1}).$ (7)
This makes intuitive sense: the velocity in point $x$ is given by all the continuations from $x$ to final points $x_{1}$, weighted by how likely that final point is given that we are at $x$. Note that to compute $u_{t}(x)$, one has to evaluate a joint integral over $D$ dimensions.
This observation leads us to propose a change in perspective. Rather than predicting the vector field directly, we can define an approximate vector field in terms of an expectation with respect to a variational distribution $q_{t}^{\theta}$ with parameters $\theta$,
$v_{t}^{\theta}(x):=\int u_{t}(x\mid x_{1})q_{t}^{\theta}(x_{1}\mid x)\mathrm{d}x_{1}.$ (8)
Clearly, in this construction $v_{t}^{\theta}(x)$ will be equal to $u_{t}(x)$ when $q_{t}^{\theta}(x_{1}\mid x)$ and $p_{t}(x_{1}\mid x)$ are identical. This implies that we can map a flow matching problem onto a variational inference problem.
Concretely, we can define a variational flow matching problem by minimizing the Kullback-Leibler (KL) divergence from $p_{t}$ to $q_{t}^{\theta}$, which we can express as (see section A.1 for derivations)
$\mathbb{E}_{t}\left[\text{KL}\big{(}p_{t}(x)p_{t}(x_{1}\mid x)~{}||~{}p_{t}(x)q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right]+\text{const},$ (9)
where $t\sim\text{Uniform}(0,1)$ and $x,x_{1}\sim p_{t}(x,x_{1})$.
This leads us to propose the variational flow matching (VFM) objective
$\mathcal{L}_{\text{VFM}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right].$ (10)
Note that, though different, this objective is similar to a typical negative log-likelihood objective.
At this stage of the exposition, it is not yet clear how useful this variational perspective on flow matching is in practice. First, while we can in principle map any flow matching problem onto a variational inference problem, this requires learning an approximation of a potentially complex, high-dimensional distribution $p_{t}(x_{1}\mid x)$. Second, we express $v^{\theta}_{t}(x)$ as an expectation, which could well be intractable. This could mean that we would have to approximate this expectation, for example using a Monte Carlo estimator. We will address these concerns in section 3.2.
3.2 Mean-Field Variational Flow Matching
Decomposing the conditional vector field.
At first glance we do not seem to obtain much from this variation view due to the intractability of $p_{t}(x_{1}\mid x)$ and $v_{t}^{\theta}(x)$. Fortunately, we can simplify the objective and the calculation of the marginal vector field under certain conditions on the conditional vector field. Specifically, we will focus on the special case where the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$, such as in straight line interpolations commonly used in flow matching. Informally, we show that in that case the specific form of the distribution used does not affect the expected value of the conditional vector field as long as the marginals coincide. As such, the solution can be found by a mean-field/factorized distribution exactly. This reduces the problem from one high-dimensional problem, into a series of low dimension problems. Formally, the following holds:
Theorem 1.
Assume that the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$. Then, for any distribution $r_{t}(x_{1}\mid x)$ such that the marginal distributions coincide with those of $p_{t}(x_{1}\mid x)$, the corresponding expectations of $u_{t}(x\mid x_{1})$ are equal, i.e.
$\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right].$ (11)
We provide a proof in section A.3. It follows directly from theorem 1 that without loss of generality we can consider the considerably easier task of a fully-factorized approximation
$q_{t}^{\theta}(x_{1}\mid x):=\prod_{d=1}^{D}q_{t}^{\theta}(x_{1}^{d}\mid x).$ (12)
We refer to this special case as mean-field variational flow matching (MF-VFM), and the VFM objective reduces to
$\mathcal{L}_{\text{MF-VFM}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right]=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x)\right].$ (13)
Computing the marginal vector field.
Now, to calculate the vector field $v_{t}^{\theta}(x)$, we can simply substitute the factorized distribution $q^{\theta}_{t}(x_{1}\mid x)$ into eq. 8. However, this still requires an evaluation of an expectation. Fortunately, leveraging the linearity condition significantly simplifies this computation. Informally, under this linearity condition, as long as we have access to the first moment of one-dimensional distributions $q^{\theta}_{t}(x_{1}^{d}\mid x)$, we can efficiently calculate $v_{t}^{\theta}(x)$. Therefore, if two distributions $q_{t}(x_{1}\mid x)$ and $q_{t}^{\prime}(x_{1}\mid x)$ share the same first moments, they will describe the same vector field, regardless of any other differences they may have. Note that the training procedure will differ for two distinct distributions – e.g. Gaussian versus Categorical – so the form of the distribution $q^{\theta}_{t}(x_{1}\mid x)$ remains practically important, a flexibility provided through the variational view on flow matching.
Formally, we can rewrite an arbitrary linear conditional vector field as
$u_{t}(x\mid x_{1})=A_{t}(x)x_{1}+b_{t}(x),$ (14)
where $A_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}\times\mathbb{R}^{D}$ and $b_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. If we substitute this into the definition of $v_{t}^{\theta}(x)$ in eq. 8, we can use the linearity of the expected value to see that
$v_{t}^{\theta}(x)=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[A_{t}(x)x_{1}+b_{t}(x)\right]=A_{t}(x)\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x).$ (15)
If we now use the standard flow matching case of using a conditional vector field based on a linear interpolation, the approximate vector field can be expressed in terms of the first moment of the variational approximation:
$v^{\theta}_{t}(x)$ $=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[\frac{x_{1}-x}{1-t}\right]=\frac{\mu_{1}-x}{1-t},$ $\mu_{1}$ $:=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[x_{1}\right].$ (16)
Note that this covers both the case of categorical data, which we focus on in this paper, and the case of continuous data, as considered in traditional flow matching methods.
At first glance, the linearity condition of the conditional vector field $u_{t}(x\mid x_{1})$ in theorem 1 might seem restricting. However, in most state-of-the-art generative modeling techniques, this condition is satisfied, e.g. diffusion-based models, such as flow matching [24, 2, 27], diffusion models [14, 46], and models that combine the injection of Gaussian noise with blurring [37, 15], among others [10, 43].
Figure 1: Reparameterized velocities by CatFlow. Note that as $\theta_{t}(x_{t})$ lies in the probability simplex, indeed all velocity implicitly are given the inductive bias of point towards it.
3.3 CatFlow: Mean-Field Variational Flow Matching for Categorical Data
The CatFlow Objective
In CatFlow, we directly apply the VFM framework to the categorical case. Let our parameterised variational distribution $q^{\theta}_{t}(x_{1}^{d}\mid x)=\mathsf{Cat}(x^{d}_{1}\mid\theta_{t}^{d}(x))$, and let us denote $\theta_{t}^{dk}(x):=q_{t}(x_{1}^{d}=k\mid x)$ for brevity. Then, the $d$th component of the learned vector field is
$v_{t}^{\theta,d}(x):=\sum_{k=1}^{K_{d}}\theta^{dk}_{t}(x)\frac{\mathbb{I}[x_{1}^{d}=k]-x}{1-t}$ (17)
Intuitively, CatFlow learns a distribution over the conditional trajectories to all corners of the probability simplices, rather than regressing towards an expected conditional trajectory.
In the categorical setting, the MF-VFM objective can be written out explicitly. Writing out the probability mass function of the categorical distribution, we see that
$\log q_{t}^{\theta}(x_{1}^{d}\mid x)=\log\prod_{k=1}^{K^{d}}(\theta_{t}^{dk}(x))^{\mathbb{I}[x_{1}^{d}=k]}=\sum_{d=1}^{D}\mathbb{I}[x_{1}^{d}=k]\log\theta_{t}^{dk}(x).$ (18)
As such, we find that CatFlow objective is given by a standard cross-entropy loss:
$\mathcal{L}_{\text{CatFlow}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\theta^{dk}_{t}(x)\right].$ (19)
Note, however, that when actually computing $v_{t}^{\theta}$, this can be done efficiently, since
$\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[u_{t}(x^{d}\mid x_{1}^{d})\right]=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[\frac{x_{1}^{d}-x^{d}}{1-t}\right]=\frac{\theta^{d}(x)-x^{d}}{1-t},$ (20)
since $\theta^{d}(x):=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}[u_{t}^{d}(x_{1}^{d}\mid x)]$ and the other terms are not in the expectation. Note that this geometrically corresponds to learning the mapping to a point in the probability simplex, and then flowing towards that. This procedure is illustrated in fig. 1. Because of this, training CatFlow is no less efficient than flow matching.
More precisely, training CatFlow offers three key benefits over standard flow matching. First, the added inductive bias ensures generative paths align with realistic trajectories, improving performance and convergence by avoiding misaligned paths. Second, using cross-entropy loss instead of mean-squared error improves gradient behavior during training, enhancing learning dynamics and speeding up convergence. Lastly, CatFlow’s ability to learn probability vectors, rather than directly choosing classes, allows it to express uncertainty about variables at a specific time. This is especially useful in complex domains like molecular generation, where initial uncertainty about components decreases as more structure is established, leading to more precise predictions.
Permutation Equivariance.
Graphs, defined by vertices and edges, lack a natural vertex order unlike other data types. This permutation invariance means any vertex labeling represents the same graph if the connections remain unchanged. Note that even though the following results apply to graphs, an unordered set of categorical variables can be described by a graph without edges. Under natural conditions – see appendix A – we ensure CatFlow adheres to this symmetry (see section A.4 for the proof).
Theorem 2.
CatFlow generates exchangeable distributions, i.e. CatFlow generates all graph permutations with equal probability.
4 Flow Matching and Score Matching: Bridging the Gap
In this section, we relate the VFM framework to existing generative modeling approaches. First, we show that VFM has standard flow matching as a special case when the variational approximation is Gaussian. This implies that VFM provides a more general approach to learning CNFs. Second, we show that through VFM, we are not only able to compute the target vector field, but also the score function as used in score-based diffusion. This has two primary theoretical implications: 1) VFM simultaneously learns deterministic and stochastic dynamics – as diffusion models rely on stochastic dynamics, and 2) VFM provides a variational bound on the model likelihood.
Relationship to Flow Matching.
Informally, the following theorem states that VFM has flow matching as a special case when the variational approximation is Gaussian and under certain assumptions on $u_{t}(x\mid x_{1})$. This result is key, as it shows that indeed optimizing a model through VFM is more flexible than regular flow matching. Formally, the following holds (see theorem 3 in section A.2 for the proof):
Theorem 3.
Assume the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$ and is of the form
$u_{t}(x|x_{1})=A_{t}(x)x_{1}+b_{t}(x),$ (21)
where $A_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}\times\mathbb{R}^{D}$ and $b_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. Moreover, assume that $A_{t}(x)$ is an invertible matrix and $q_{t}^{\theta}(x_{1}\mid x)=\mathcal{N}(x_{1}\mid\mu_{t}^{\theta}(x),\Sigma_{t}(x))$, where $\Sigma_{t}(x)=\frac{1}{2}(A^{\top}_{t}(x)A_{t}(x))^{-1}$. Then, VFM reduces to flow matching.
Relationship to Score-Based Models.
Flow matching [24] is inspired by score-based models [46] and shares strong connections with them. In score-based models, the objective is to approximate the score function $\nabla_{x}\log p_{t}(x)$ with a function $s^{\theta}_{t}(x)$. A connection to VFM becomes apparent by observing that the score function can also be expressed as an expectation with respect to $p_{t}(x_{1}\mid x)$ (see section A.5 for derivation):
$\nabla_{x}\log p_{t}(x)$ $=\int p_{t}(x_{1}\mid x)\nabla_{x}\log p_{t}(x\mid x_{1})\mathrm{d}x_{1}=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ (22)
where $\nabla_{x}\log p_{t}(x\mid x_{1})$ is the tractable conditional score function. Similarly, we can parameterize $s^{\theta}_{t}(x)$ in terms of an expectation with respect to a variational approximation $q^{\theta}_{t}(x_{1}\mid x)$,
$s_{t}^{\theta}(x):=\int q_{t}^{\theta}(x_{1}\mid x)\nabla_{x}\log p_{t}(x\mid x_{1})\>\mathrm{d}x_{1}.$ (23)
It is now clear that $s_{t}^{\theta}(x)=\nabla_{x}\log p_{t}(x)$ when $q_{t}^{\theta}(x_{1}\mid x)=p_{t}(x_{1}\mid x)$. This suggests that there exists a variational formulation of score-based models that is entirely analogous to VFM. More specifically, we see that both models could be optimized using a single objective, which yields an approximation $q_{t}^{\theta}(x_{1}\mid x)$ that parameterizes both $v_{t}^{\theta}(x)$ and $s^{\theta}_{t}(x)$.
Following [46, 3], we can construct stochastic generative dynamics $dx=\tilde{v}_{t}^{\theta}(x)dt+g_{t}dw$ to approximate the true dynamics $dx=\tilde{u}_{t}(x)dt+g_{t}dw$ (see details in section A.6), with
$\tilde{u}_{t}(x)$ $:=\mathbb{E}_{p_{t}(x_{1}|x)}\left[u_{t}(x\mid x_{1})+\frac{g_{t}^{2}}{2}\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ $\tilde{v}^{\theta}_{t}(x)$ $:=v_{t}^{\theta}(x)+\frac{g_{t}^{2}}{2}s_{t}^{\theta}(x).$ (24)
Here $g_{t}:[0,1]\rightarrow\mathbb{R}_{+}$ is a scalar function, and $w$ is a standard Wiener process.
This connection has two important implications. First, it shows that learning a variational approximation $q^{\theta}_{t}(x_{1}\mid x)$ can be used to define both deterministic and stochastic dynamics, whereas flow matching typically considers deterministic dynamics only (as flows are viewed through the lens of ODEs). Second, it enables us to show that a reweighted version of the VFM objective provides a bound on the log-likelihood of the model. This bound provides another theoretical motivation for learning using the VFM objective.
Theorem 4.
Rewrite the Variational Flow Matching objective as follows:
$\mathcal{L}_{\text{VFM}}(\theta)=\mathbb{E}_{t,x}\left[\mathcal{L}^{\theta}(t,x)\right]\quad\textrm{where}\quad\mathcal{L}^{\theta}(t,x)=-\mathbb{E}_{x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right].$ (25)
Then, the following holds:
$-\mathbb{E}_{x_{1}}\left[\log q_{1}^{\theta}(x_{1})\right]\leq\mathbb{E}_{t,x}\left[\lambda_{t}(x)\mathcal{L}^{\theta}(t,x)\right]+C,$ (26)
where $\lambda_{t}(x)$ is a non-negative function and $C$ is a constant.
We provide a proof in section A.7. We further note that that applying the same linearity condition that we discussed in section 3.2 to the conditional score function maintains all the same connections with score-based models.
5 Related Work
Diffusion Models for Discrete Data.
Several approaches to discrete generation using diffusion models have been developed [4, 28, 53, 17]. For graph generation specifically, [50] utilize a Markov process that progressively edits graphs by adding or removing edges and altering node or edge categories and is trained using a graph transformer network, that reverses this process to predict the original graph structure from its noisy version. This approach breaks down the complex task of graph distribution learning into simpler node and edge classification tasks. Moreover, [19] proposes a score-based generative model for graph generation using a system of stochastic differential equations (SDEs). The model effectively captures the complex dependencies between graph nodes and edges by diffusing both node features and adjacency matrices through continuous-time processes. These are non-autoregressive graph generation approaches that perform on par with autoregressive ones, such as in [26, 23, 31]. Other non-autoregressive approaches worth mentioning are [25, 29, 28].
Flow-based methods for Discrete Data.
Very recently two flow-based methods for discrete generative modeling have been proposed, which differ both in terms of technical approach and intended use case from the work that we present here. [^2] Flow matching and diffusion models have also been proposed for geometric graph generation, e.g. in [20, 47] and [16, 49] respectively, but since these approaches are continuous (as they generate coordinates based on some conformer) they consider a fundamentally different task than the ones we consider here.
In [48], a Dirichlet flow framework for DNA sequence design is introduced, utilizing a transport problem defined over the probability simplex. This approach differs from CatFlow in that it represents the conditional probability path $p_{t}(x\mid x_{1})$ using a Dirichlet distribution. This implies that points $x$ are constrained to the simplex, which is not the case for CatFlow. Dirichlet Flows have not been evaluated on graph generation, but we did carry out preliminary experiments based on released source code. We have opted not to report these results; we did not obtain good performance out of the box, but have also not invested substantial time in architecture and hyperparameter selection.
In [6], Discrete Flow Models (DFMs) are introduced. DFMs use Continuous-Time Markov Chains to enable flexible and dynamic sampling in multimodal generative modeling of both continuous and discrete data. Though sharing a goal, this approach differs significantly from CatFlow as in the end the resulting model does not learn a CNF, but rather generation through sequential sampling from a time-dependent categorical distribution. As in the case of Dirichlet flows, no evaluation on graph generation was performed.
The switch to the variational perspective is inspired by [50], showing significant improvement through viewing the dynamics as a classification task over end points. However, CatFlow is still a continuous model, and integrates – rather than iteratively samples – during generation.
Table 1: Results abstract graph generation.
Ego-small Community-small
Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
GraphVAE 0.130 0.170 0.050 0.350 0.980 0.540
GNF 0.030 0.100 0.001 0.200 0.200 0.110
EDP-GNN 0.052 0.093 0.007 0.053 0.144 0.026
GDSS 0.021 0.024 0.007 0.045 0.086 0.007
CatFlow 0.013 0.024 0.008 0.018 0.086 0.007
6 Experiments
We evaluate CatFlow in three sets of experiments. First, we consider an abstract graph generation task proposed in [30], where the goal of this task is to evaluate if CatFlow is able to capture the topological properties of graphs. Second, we consider two common molecular benchmarks, QM9 [34] and ZINC250k [18], consisting of small and (relatively) large molecules respectively. This task is chosen to see if CatFlow can learn semantic information in graph generation, such as molecular properties. Finally, we perform an ablation comparing CatFlow to standard flow matching, specifically in terms of generalization. The experimental setup and model choices are provided in appendix C.
Note that we treat graphs as purely categorical/discrete objects and do not consider ‘geometric’ graphs that are embedded in e.g. Euclidean space. Specifically, for some graph with $K_{v}$ node classes and $K_{e}$ edge classes, we process the graph as a fully-connected graph, where each node is treated as a categorical variable of one of $K_{v}$ classes and each edge of $K_{e}+1$ classes, where the extra class corresponds with being absent.
Table 2: Results molecular generation.
QM9 ZINC250k
Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
MoFlow 91.36 98.65   4.467 63.11 99.99 20.931
EDP-GNN 47.52 99.25   2.680 82.97 99.79 16.737
GraphEBM   8.22 97.90   6.143   5.29 98.79 35.471
GDSS 95.72 98.46   2.900 97.01 99.64 14.656
Digress 99.00 96.20 - - - -
CatFlow 99.81 99.95   0.441 99.21 100.00 13.211
Figure 2: CatFlow samples of QM9 (top) and ZINC250k (bottom).
6.1 Abstract graph generation
We first evaluate CatFlow on an abstract graph generation task, including synthetic and real-world graphs. We consider 1) Ego-small (200 graphs), consisting of small ego graphs drawn from a larger Citeseer network dataset [41], 2) Community-small (100 graphs), consisting of randomly generated community graphs, 3) Enzymes (587 graphs), consisting of protein graphs representing tertiary structures of the enzymes from [40], and 4) Grid (100 graphs), consisting of 2D grid graphs. We follow the standard experimental setup popularized by [54] and hence report the maximum mean discrepancy (MMD) to compare the distributions of degree, clustering coefficient, and the number of occurrences of orbits with 4 nodes between generated graphs and a test set. Following [19], we also use the Gaussian Earth Mover’s Distance kernel to compute the MMDs instead of the total variation.
The results of the Ego-small and Community-small tasks are summarized in table 1, additional results (and error bars) are provided in appendix B. The results indicate that CatFlow is able to capture topological properties of graphs, and performs well on abstract graph generation.
6.2 Molecular Generation: QM9 & ZINC250k
Molecular generation entails designing novel molecules with specific properties, a complex task hindered by the vast chemical space and long-range dependencies in molecular structures. We evaluate CatFlow on two popular molecular generation benchmarks: QM9 and ZINC250k [34, 18].
We follow the standard setup – e.g. as in [42, 28, 50, 19] – of kekulizing the molecules using RDKit [22] and removing the hydrogen atoms. We sample 10,000 molecules and evaluate them on validity, uniqueness, and Fréchet ChemNet Distance (FCD) – evaluating the distance between data and generated molecules using the activations of ChemNet [33]. Here, validity is computed without valency correction or edge resampling, hence following [55] rather than [42, 28], as is more reasonable due to the existence of formal charges in the data itself. We do not report novelty for QM9 and ZINC250k, as QM9 is an exhaustive list of all small molecules under some chemical constraint and all models obtain (close to) 100% novelty on ZINC250k. [^3] CatfFlow obtains 49% novelty on QM9.
The results are summarized in table 2 and samples from the model are shown in fig. 2. CatFlow obtains state-of-the-art results on both QM9 and ZINC250k, virtually obtaining perfect performance on both datasets. It is worth noting that CatFlow also converges faster than flow matching and is not computationally more expensive than any of the baselines either during training or generation.
6.3 CatFlow Ablations
(a) Score $100\%$ of data.
(b) Score $20\%$ of data.
(c) Score $5\%$ of data.
Figure 3: Ablation results, comparison between standard flow matching and CatFlow.
To understand the difference in performance between CatFlow and a standard flow matching formulation we perform ablations. we focus on generalization capabilities, and as such consider ablations that the number of parameters in the model and the amount of training data.
In fig. 3 we report a score, which is the percentage of generated molecules that is valid and unique. CatFlow not only outperforms regular flow matching in the large model and full data setting, but is also significantly more robust to a decrease in model-size and data. Moreover, we observe significantly faster convergence (curves not shown). We hypothesize this is a consequence of the optimization procedure not exploring ‘irrelevant’ paths that do not point towards the probability simplex.
7 Conclusion
We have introduced a variational reformulation of flow matching. This formulation in turn informed the design of a simple flow matching method for categorical data, which achieves strong performance on graph generation tasks. Variational flow is very general and opens up several lines of inquiry. We see immediate opportunities to apply CatFlow to other discrete data types, including text, source code, and more broadly to the modeling of mixed discrete-continuous data modalities. Additionally, the connections to score-based models that we identify in this paper, suggest a path towards learning both deterministic and stochastic dynamics.
Limitations.
While the VFM formulation that we identify in this paper has potential in terms of its generality, we have as yet only considered its application to the specific task of categorical graph generation. We leave other use cases of VFM to future work. A limitation of CatFlow, which is shared with related approaches to graph generation, is that reasoning about the set of possible edges has a cost that is quadratic in the number of nodes. As a result CatFlow does not scale well to e.g. large proteins of $10^{4}$ or more atoms.
Ethics Statement.
Graph generation in general, and molecular generation specifically, holds great promise for advancing drug discovery and personalized medicine. However, this technology also poses ethical concerns, such as the potential for misuse in creating harmful substances. In terms of technology readiness, this work is not yet at a level where we foresee direct benefits or risks.
Acknowledgements
This project was supported by the Bosch Center for Artificial Intelligence.
Appendix A Proofs and Derivations
A.1 Derivation of the Variational Flow Matching Objective
We derive eq. 9 that states the equivalence of the optimization of the VFM objective and minimization of the KL divergence between the true endpoint distribution $p_{t}(x_{1}\mid x)$ and the variational approximation $q^{\theta}_{t}(x_{1}\mid x)$. Note that
$\min_{\theta}~{}\mathbb{E}_{t,x}\left[\text{KL}\big{(}p_{t}(x_{1}\mid x)~{}||~{}q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=\max_{\theta}~{}\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right],$ (27)
where $t\sim\text{Uniform}(0,1)$, $x\sim p_{t}(x)$ and $x_{1}\sim p_{\text{data}}(x_{1})$.
First, we rewrite the KL divergence as combination of entropy and cross-entropy $\text{KL}\left(p~{}||~{}q\right)=\text{H}(p,q)-\text{H}(p)$:
$\mathbb{E}_{t,x}\left[\text{KL}\big{(}p_{t}(x_{1}\mid x)~{}||~{}q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=$ $\mathbb{E}_{t,x}\left[\text{H}\big{(}p_{t}(x_{1}\mid x),q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]-$ (28)
$\mathbb{E}_{t,x}\left[\text{H}\big{(}p_{t}(x_{1}\mid x)\big{)}\right].$ (29)
We observe that the entropy term does not depend on the parameters $\theta$. Consequently we disregard it when optimising the variational distribution $q^{\theta}_{t}(x_{1}\mid x)$.
Second, we rewrite the cross-entropy term:
$\mathbb{E}_{t,x}\left[\text{H}\big{(}p_{t}(x_{1}\mid x),q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right].$ (30)
Therefore, the second part of eq. 27 corresponds to negative cross-entropy and minimisation corresponds to maximisation of negative cross-entropy.
A.2 Flow Matching as a Special Case of Variational Flow Matching
See 3
Proof.
Let us substitute the assumed form of $q_{t}^{\theta}(x_{1}\mid x)$ into the VFM objective:
$\mathcal{L}_{\text{VFM}}(\theta)$ $=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right]$ (31)
$=-\mathbb{E}_{t,x,x_{1}}\left[\log\left(\left(2\pi\right)^{-D/2}\left|\Sigma_{t}(x)\right|^{-1/2}\exp\left(-\left\|A_{t}(x)\left(x_{1}-\mu_{t}^{\theta}(x)\right)\right\|_{2}^{2}\right)\right)\right]$ (32)
$=\mathbb{E}_{t,x,x_{1}}\left[\left\|A_{t}(x)\left(x_{1}-\mu_{t}^{\theta}(x)\right)\right\|_{2}^{2}\right]+\frac{1}{2}\mathbb{E}_{t,x,x_{1}}\Big{[}D\log\left(2\pi\right)+\log\left|\Sigma_{t}(x)\right|\Big{]}$ (33)
$=\mathbb{E}_{t,x,x_{1}}\left[\left\|\left(A_{t}(x)x_{1}+b_{t}(x)\right)-\left(A_{t}(x)\mu_{t}^{\theta}(x)+b_{t}(x)\right)\right\|_{2}^{2}\right]+C$ (34)
$=\mathbb{E}_{t,x,x_{1}}\left[\left\|u_{t}(x\mid x_{1})-v_{t}^{\theta}(x)\right\|_{2}^{2}\right]+,$ (35)
which is what we wanted to show. ∎
A.3 Decomposition of the Flow
See 1
Proof.
Applying the linearity condition, we rewrite the conditional vector field $u_{t}(x|x_{1})$ as such:
$u_{t}(x\mid x_{1})=A_{t}(x)x_{1}+b_{t}(x),$ (36)
where $A_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}\times\mathbb{R}^{D}$ and $b_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$. Then, we substitute it into equation:
$\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]$ $=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[A_{t}(x)x_{1}+b_{t}(x)\right]$ (37)
$=A_{t}(x)\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x).$ (38)
Then, we know, that for any distribution $r$ such that the marginal distributions coincide with those of $p$ the following holds:
$\mathbb{E}_{r}[x]=\mathbb{E}_{p}[x].$ (39)
Applying this fact to eq. 38 we obtain:
$\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]$ $=A_{t}(x)\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x)$ (40)
$=\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[A_{t}(x)x_{1}+b_{t}(x)\right]$ (41)
$=\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right],$ (42)
which is what we wanted to show.∎
A.4 CatFlow
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be a graph with node and edge features given by $\mathbf{H}_{n}\in\mathbb{R}^{|\mathcal{V}|\times d_{n}}$ and $\mathbf{H}_{e}\in\mathbb{R}^{|\mathcal{V}|\times|\mathcal{V}|\times d_{e}}$ respectively, and let $x$ denote the graphs and its features. Moreover, let $\pi\in S_{|\mathcal{V}|}$ be a permutation and $\mathbf{P}$ its associated permutation matrix, such that the action of the group is defined as
  • $\pi\cdot\mathbf{H}_{n}:=\mathbf{P}\mathbf{H}_{n}$,
  • $\pi\cdot\mathbf{H}_{e}:=\mathbf{P}\mathbf{H}_{e}\mathbf{P}^{\top}$.
To simplify notation, we will simply write $\pi\cdot x$ to denote the above operation.
Lemma 1.
If $\theta_{t}(x)$ is permutation equivariant w.r.t $S_{|\mathcal{V}|}$, then so is $v_{t}$.
Proof.
Note that
$v_{t}(\pi\cdot x)=\frac{\theta_{t}(\pi\cdot x)-\pi\cdot x}{1-t}=\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t},$
where the last step follows from permutation equivariance. Moreover, since $\pi$ acts on $x$ through permutation matrices, we can leverage the distributive property of linear operators, i.e. we conclude that
$\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t}=\pi\cdot\frac{(\theta_{t}(x)-x)}{1-t}=\pi\cdot v_{t}(x),$
finishing the proof.∎
Theorem 5.
Let $p_{0}$ be an exchangeable distribution – e.g. a standard normal distribution – and $\theta_{t}(x)$ be permutation equivariant. Then, all permutations of graphs are generated with equal probability.
Proof.
[…]
where again the last step follows by basic properties of linear operators. Therefore, since $p_{0}$ assigns equal density of all permutations of $x$, the resulting distribution $p_{1}$ preservers this property, which is what we wanted to show.∎
A.5 Derivation of the Score Function
In this subsection we want to derive the equation that allows us to express the score function $\nabla_{x}\log p_{t}(x)$ in terms of the posterior probability path $p_{t}(x\mid x_{1})$. Notice that
$\nabla_{x}\log p_{t}(x)$ $=\frac{1}{p_{t}(x)}\nabla_{x}p_{t}(x)$ (43)
[…]
$=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right].$ (48)
A.6 Stochastic Dynamics with Variational Flow Matching
In this subsection, we discuss how the VFM framework can be applied to construct stochastic generative dynamics and how it relates to score-based models.
First, let us consider the marginal vector field $u_{t}(x)$. It provides the deterministic dynamic that can be written as the following ordinary differential equation (ODE):
$dx=u_{t}(x)dt.$ (49)
For the vector field $u_{t}(x)$ we know that – starting from the distribution $p_{0}(x)$ – it generates some probability path $p_{t}(x)$. However, as we know from [46, 3], if we have access to the score function $\nabla_{x}\log p_{t}(x)$ of distribution $p_{t}(x)$, we can construct a stochastic differential equation (SDE) that, staring from distribution $p_{0}(x)$, generates the same probability path $p_{t}(x)$:
$dx=\left[u_{t}(x)+\frac{g_{t}^{2}}{2}\nabla_{x}\log p_{t}(x)\right]dt+g_{t}dw,$ (50)
where $g:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is a scalar function, and $w$ is a standard Wiener process.
Given that
$u_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]\quad\textrm{and}\quad\nabla_{x}\log p_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ (51)
we can rewrite eq. 50 in the following form:
$dx=\tilde{u}_{t}(x)dt+g_{t}dw,\quad\textrm{where}$ (52)
$\tilde{u}_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\tilde{u}_{t}(x\mid x_{1})\right]$ $\quad\textrm{and}\quad\tilde{u}_{t}(x\mid x_{1})=u_{t}(x\mid x_{1})+\frac{g_{t}^{2}}{2}\nabla_{x}\log p_{t}(x\mid x_{1}).$ (53)
Importantly, the function $g_{t}$ does not affect the distribution path $p_{t}(x)$: it only changes the stochasticity of the trajectories. In the extreme case when $g_{t}\equiv 0$, the SDE in eq. 50 coincides with the ODE in eq. 49.
Moreover, if we construct the stochastic dynamics in this way, we obtain score-based models as a special case. In score-based models [46] the deterministic process that corresponds to probability path $p_{t}(x)$ has the following form:
$dx=u_{t}(x)dt\quad\textrm{where}\quad u_{t}(x)=f_{t}(x)+\frac{g_{t}^{2}}{2}\nabla_{x}\log p_{t}(x),$ (54)
where $f_{t}(x)$ is tractable. Substituting this $u_{t}(x)$ into eq. 50 we obtain:
$dx=\left[f_{t}(x)+g_{t}^{2}\nabla_{x}\log p_{t}(x)\right]dt+g_{t}dw.$ (55)
The SDE in eq. 55 only depends on the score function, so in score-based models, the aim is to learn the score function.
Now, let us note that similarly to vector field $u_{t}(x)$, the drift term $\tilde{u}_{t}(x)$ in eq. 52 can be expressed in terms of an end point distribution $p_{t}(x_{1}\mid x)$. Consequently, having a variational approximation of end point distribution $q_{t}^{\theta}(x_{1}\mid x)$, allows us to construct an approximation of the drift term $\tilde{u}_{t}(x)$:
$\tilde{v}^{\theta}_{t}(x)=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x)}\left[\tilde{u}_{t}(x\mid x_{1})\right]$ $=v^{\theta}_{t}(x)+\frac{g_{t}^{2}}{2}s^{\theta}_{t}(x),$ (56)
$\textrm{where}\quad v^{\theta}_{t}(x)=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]\quad\textrm{and}\quad$ $s^{\theta}_{t}(x)=\mathbb{E}_{q_{t}^{\theta}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right].$ (57)
Thus, we may define the following approximated SDE:
$dx=\tilde{v}^{\theta}_{t}(x)dt+g_{t}dw\quad\textrm{or}\quad dx=\left[v^{\theta}_{t}(x)+\frac{g_{t}^{2}}{2}s^{\theta}_{t}(x)\right]dt+g_{t}dw.$ (58)
Then, eq. 58 is not just simply a new dynamic, it is a family of dynamics that admits deterministic dynamics as a special case when $g_{t}\equiv 0$. Importantly, the only thing we need to construct the stochastic process in eq. 58 is an approximation of the end point distributions $q_{t}^{\theta}(x_{1}\mid x)$. Additionally we know that when $p_{t}(x_{1}\mid x)=q_{t}^{\theta}(x_{1}\mid x)$, $\tilde{u}_{t}(x)=\tilde{v}^{\theta}_{t}(x)$, the processes coincide for all functions $g_{t}$. Therefore, we can train the model with the same objective as in VFM.
A.7 Variational Flow Matching as a Variational Bound on the Log-likelihood
In this subsection, we leverage the connections of VFM with stochastic processes to show that a reweighted integral over the point-wise VFM objective defines a bound on the data likelihood in the generative model.
See 4
Proof.
Let us consider the two stochastic processes, that we discussed in section A.6:
$dx=\tilde{u}_{t}(x)dt+g_{t}dw,\quad\textrm{where}\quad dx=\tilde{v}^{\theta}_{t}(x)dt+g_{t}dw.$ (59)
Note that they both start from the same prior distribution $p_{0}(x)$. The first one, by design, generates probability path $p_{t}(x)$ and ends up in the data distribution $p_{\text{data}}(x)=p_{1}(x)$. The second process generates some probability path $q^{\theta}_{t}(x)$ that depends on the variational distribution $q^{\theta}_{t}(x_{1}|x)$.
We want to find a variational bound on KL divergence between $p_{1}(x)$ and $q^{\theta}_{1}(x)$. We start by applying the result from [1] (see Lemma 2.22):
$\text{KL}\left(p_{1}(x_{1})\|q^{\theta}_{1}(x_{1})\right)$ $\leq\mathbb{E}_{t,x}\left[\frac{1}{2g_{t}^{2}}\left\|\tilde{u}_{t}(x)-\tilde{v}^{\theta}_{t}(x)\right\|_{2}^{2}\right]$ (60)
$=\mathbb{E}_{t,x}\left[\frac{1}{2g_{t}^{2}}\left\|\int\left(p_{t}(x_{1}\mid x)-q^{\theta}_{t}(x_{1}\mid x)\right)\tilde{u}_{t}(x|x_{1})dx_{1}\right\|_{2}^{2}\right]$ (61)
$\leq\mathbb{E}_{t,x}\left[\frac{1}{2g_{t}^{2}}\left(\int\Big{\|}\left(p_{t}(x_{1}\mid x)-q^{\theta}_{t}(x_{1}\mid x)\right)\tilde{u}_{t}(x|x_{1})\Big{\|}dx_{1}\right)^{2}\right]$ (62)
$\leq\mathbb{E}_{t,x}\left[\frac{1}{2g_{t}^{2}}\left(\int\left|p_{t}(x_{1}\mid x)-q^{\theta}_{t}(x_{1}\mid x)\right|\left\|\tilde{u}_{t}(x|x_{1})\right\|dx_{1}\right)^{2}\right].$ (63)
Now, let us introduce two auxiliary functions:
$l_{t}(x)=\sup_{x_{1}}\left\|\tilde{u}_{t}(x|x_{1})\right\|\quad\textrm{and}\quad\lambda_{t}(x)=\frac{l_{t}(x)}{g_{t}^{2}}.$ (64)
If we utilise $\lambda_{t}(x)$ to write down the following bound, we see that
$\text{KL}\left(p_{1}(x_{1})\|q^{\theta}_{1}(x_{1})\right)$ $\leq\mathbb{E}_{t,x}\left[\frac{\lambda_{t}(x)}{2}\left(\int\left|p_{t}(x_{1}\mid x)-q^{\theta}_{t}(x_{1}\mid x)\right|dx_{1}\right)_{2}^{2}\right].$ (65)
Next, we can apply Pinsker’s inequality, which states that for two probability distributions $p$ and $q$ the following holds:
[…]
Applying it to the inner integral, we have:
$\text{KL}\left(p_{1}(x_{1})\|q^{\theta}_{1}(x_{1})\right)$ $\leq\mathbb{E}_{t,x}\left[\lambda_{t}(x)\text{KL}\Big{(}p_{t}(x_{1}\mid x)\|q^{\theta}_{t}(x_{1}\mid x)\Big{)}\right].$ (67)
We may rewrite the left part of inequality as a combination of the data entropy and the model’s likelihood, where only the likelihood depends on parameters $\theta$:
$\text{KL}\left(p_{1}(x_{1})\|q^{\theta}_{1}(x_{1})\right)=-\text{H}\left(p_{1}(x_{1})\right)-\mathbb{E}_{x_{1}}\left[\log q^{\theta}_{1}(x_{1})\right].$ (68)
The right part of inequality can be rewritten as an expectation of entropy that does not depend on any parameters $\theta$ plus a reweighted VFM objective with weighting coefficient $\lambda_{t}(x)$:
$\mathbb{E}_{t,x}\left[\lambda_{t}(x)\text{KL}\Big{(}p_{t}(x_{1}\mid x)\|q^{\theta}_{t}(x_{1}\mid x)\Big{)}\right]=$ $-\mathbb{E}_{t,x}\left[\lambda_{t}(x)\text{H}(p_{t}(x_{1}\mid x))\right]$ (69)
$-\mathbb{E}_{t,x,x_{1}}\left[\lambda_{t}(x)q^{\theta}_{t}(x_{1}\mid x)\right]$ (70)
We see that this reweighted version of the VFM objective defines an upper bound on the model likelihood, which was what we wanted to show.∎
A.8 Stochastic Dynamics under Linearity Conditions
In this subsection, we discuss the connection between VFM and stochastic dynamics under the condition of linearity in $x_{1}$ of the conditional vector field $u_{t}(x\mid x_{1})$ and conditional score function $\nabla_{x}\log p_{t}(x\mid x_{1})$.
As we discuss in section 3.2, under the linearity condition, we may express $u_{t}(x)$ in terms of any distribution of end points $r_{t}(x_{1}\mid x)$ if it has the same marginals as $p_{t}(x_{1}\mid x)$. In section A.5, we demonstrate that the score function $\nabla_{x}\log p_{t}(x)$ can also be expressed in terms of end point distributions $p_{t}(x_{1}\mid x)$. Therefore, the score function may also be equally expressed in terms of distribution $r_{t}(x_{1}\mid x)$ if it has the same marginals as $p_{t}(x_{1}\mid x)$. This fact is easy to show in the same way as we present in section A.3.
Consequently, under the linearity conditions, the drift term $\tilde{u}_{t}(x)$ can also be expressed in terms of $r_{t}(x_{1}\mid x)$, as it is just a linear combination of the vector field $u_{t}(x)$ and score function $\nabla_{x}\log p_{t}(x)$. Hence, the transition from the distribution $p_{(}x_{1}\mid x)$ to some distribution $r_{(}x_{1}\mid x)$ does not affect the discussion of connections of stochastic dynamics in section A.6.
Furthermore, the transition from distribution $p_{t}(x_{1}\mid x)$ to some distribution $r_{t}(x_{1}\mid x)$ does not affect connections of the VFM objective with the model likelihood that we discuss in section A.7. It is easy to see that in the derivations, we only rely on functions $\tilde{u}_{t}(x)$ and $\tilde{v}^{\theta}_{t}(x)$ (see eq. 60). However, as we discussed, they are not affected by the transition from $p_{t}(x_{1}\mid x)$ to some $r_{t}(x_{1}\mid x)$. Therefore, we may repeat all the same derivations for some $r_{(}x_{1}|x)$ using a factorized distribution.
Appendix B Detailed Results
Here, we provide the results for CatFlow with standard deviations, as computed as in [19] through multiple seeds.
Ego-small Community-small
$4\leq|\mathcal{V}|\leq 18$ $12\leq|\mathcal{V}|\leq 20$
Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
$0.013\pm 0.007$ $0.024\pm 0.009$ $0.008\pm 0.005$ $0.018\pm 0.012$ $0.086\pm 0.021$ $0.007\pm 0.005$
Enzymes Grid
$10\leq|\mathcal{V}|\leq 125$ $100\leq|\mathcal{V}|\leq 400$
Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
$0.013\pm 0.012$ $0.062\pm 0.011$ $0.008\pm 0.007$ $0.115\pm 0.010$ $0.004\pm 0.002$ $0.075\pm 0.071$
QM9 ZINC250k
$1\leq|\mathcal{V}|\leq 9$, 4 atom types $6\leq|\mathcal{V}|\leq 38$, 9 atom types
Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
$99.81\pm 0.03$ $99.95\pm 0.02$ $0.441\pm 0.023$ $99.21\pm 0.04$ $100.00\pm 0.00$ $13.211\pm 0.12$
Appendix C Experimental setup
C.1 Model
To ensure a comparison on equal terms to baselines, we employ the same graph transformer network as proposed in [11], which was also used in [50], along with the same hyper-parameter setup. We summarize the parametrization of our network here.
Just as done in DiGress, our graph transformer takes as input a graph $(\mathbf{H}_{n},\mathbf{H}_{e})$ and predicts a distribution over the clean graphs, using structural and spectral features to improve the network expressivity, which we denote as $\mathbf{H}_{g}$. Each transformer layer does the following operations:
  1. 1.
    Node Features $\mathbf{H}_{n}$ and Edge Features $\mathbf{H}_{e}$:
    1. (a)
      Linear Transformation: Apply linear transformations to both $\mathbf{H}_{n}$ and $\mathbf{H}_{e}$.
    2. (b)
      Outer Product and Scaling: Compute the outer product of the transformed features and apply scaling.
  2. 2.
    Node Features $\mathbf{H}_{n}$:
    1. (a)
      Feature-wise Linear Modulation (FiLM): Apply FiLM to the transformed node features using global features $\mathbf{H}_{g}$.
  3. 3.
    Edge Features $\mathbf{H}_{e}$:
    1. (a)
      Feature-wise Linear Modulation (FiLM): Apply FiLM to the transformed edge features using global features $\mathbf{H}_{g}$.
  4. 4.
    Self-Attention Mechanism:
    1. (a)
      Linear Transformation: Apply a linear transformation to the transformed node features.
    2. (b)
      Softmax Operation: Compute the attention scores using the softmax function.
    3. (c)
      Attention Score Calculation: Calculate the weighted sum of the transformed node features based on the attention scores.
  5. 5.
    Global Features $y$:
    1. (a)
      Pooling: Apply PNA pooling to the node features $\mathbf{H}_{n}$ and edge features $\mathbf{H}_{e}$.
    2. (b)
      Summation: Sum the pooled features with the global features $\mathbf{H}_{g}$.
  6. 6.
    Final Outputs:
    1. (a)
      Node Features $\mathbf{H}_{n}^{\prime}$: Obtain updated node features after the attention mechanism.
    2. (b)
      Edge Features $\mathbf{H}_{e}^{\prime}$: Obtain updated edge features after the attention mechanism.
    3. (c)
      Global Features $\mathbf{H}_{g}^{\prime}$: Obtain updated global features after summation.
Here, the FiLM operation is defined as:
$\text{FiLM}(M_{1},M_{2})=M_{1}W_{1}+(M_{1}W_{2})\odot M_{2}+M_{2}$
for learnable weight matrices $W_{1}$ and $W_{2}$, and PNA is defined as:
$\text{PNA}(X)=\text{cat}(\text{max}(X),\text{min}(X),\text{mean}(X),\text{std}(X))W.$
C.2 Hyperparameters and Computational Costs
We report the hyperparameters here:
Hyperparameter Abstract QM9/ZINC250k Ablation
Optimizer AdamW AdamW AdamW
Scheduler Cosine Annealing Cosine Annealing Cosine Annealing
Learning Rate $2\cdot 10^{-4}$ $2\cdot 10^{-4}$ $2\cdot 10^{-4}$
Weight Decay $1\cdot 10^{-12}$ $1\cdot 10^{-12}$ $1\cdot 10^{-12}$
EMA 0.999 0.999 0.999
Table 3: Hyperparameter setup.
All models were trained until convergence. Furthermore, all data splits are kept the same as in [19], and hidden dimensions are kept the same as [50]. All experiments were run on a single NVIDIA RTX 6000 and took about a day to run.
Appendix D Detail CNFs
To compute the resulting distribution $p_{t}$ for CNF, one can use the change of variables formula:
$[\varphi_{t}]_{*}p_{0}(x)=p_{0}(\varphi_{t}^{-1}(x))\text{det}\left[\frac{\partial\varphi_{t}^{-1}}{\partial x}(x)\right].$ (71)
[…]
One way to assess if a vector field generates a specific probability path is using the continuity equation, i.e. we can assess whether $v_{t}$ and $p_{t}$ satisfy
$\frac{\partial}{\partial t}p_{t}(x)+\nabla\cdot(p_{t}(x)v_{t}(x))=0,$ (72)
where $\nabla$ is the divergence operator. Note that by sampling $x_{0}\sim p_{0}$, a new sample from $p_{1}$ can be generated by following this ODE, i.e. integrating
$x_{1}=x_{0}+\int_{0}^{1}v_{t}(x_{t})\mathrm{d}t.$ (73)

### v2

@@
 CatFlow is easy to implement, computationally efficient, and achieves strong results on graph generation tasks.
-In VFM, the objective is to approximate the posterior probability path, which is a distribution over possible end points of a trajectory.
-We show that VFM admits both the CatFlow objective and the original flow matching objective as special cases.
+The key observation in VFM is that we can parameterize the vector field of the flow in terms of a variational approximation of the posterior probability path, which is the distribution over possible end points of a trajectory.
+We show that this variational interpretation admits both the CatFlow objective and the original flow matching objective as special cases.
 We also relate VFM to score-based models, in which the dynamics are stochastic rather than deterministic, and derive a bound on the model likelihood based on a reweighted VFM objective.
 We evaluate CatFlow on one abstract graph generation task and two molecular generation tasks.
-In all cases, CatFlow exceeds or matches performance of the current state-of-the-art models.
+In all cases, CatFlow exceeds or matches performance of the current state-of-the-art.
 **footnotetext: These authors contributed equally to this work.
@@
 In recent years, the field of generative modeling has seen notable advancements.
-In image generation [35, 38], the development and refinement of diffusion-based approaches — specifically those using denoising score matching [51] — have proven effective for generation at scale [14, 46].
-However, while training can be done effectively, the constrained space of sampling probability paths in a diffusion requires tailored techniques to work [44, 56].
-This is in contrast to more flexible approaches such as continuous normalizing flows (CNFs) [8], that are able to learn a more general set of probability paths than diffusion models [45], at the expense of being expensive to train as they require one to solve an ODE during each training step (see e.g.
-[5, 39, 13]).
-Recently, Lipman and collaborators [24] proposed flow matching (FM), an efficient and simulation-free approach to training CNFs.
+In image generation ramesh2022hierarchical ; rombach2022high , the development and refinement of diffusion-based approaches — specifically those using denoising score matching vincent2011connection — have proven effective for generation at scale ho2020denoising ; song2020score .
+However, while training can be done effectively, the constrained space of sampling probability paths in a diffusion requires tailored techniques to work song2020denoising ; zhang2022fast .
+This is in contrast to more flexible approaches such as continuous normalizing flows (CNFs) chen2018neural , that are able to learn a more general set of probability paths than diffusion models song2021maximum , at the expense of being expensive to train as they require one to solve an ODE during each training step (see e.g.
+ben2022matching ; rozen2021moser ; grathwohl2018scalable ).
+Recently, Lipman and collaborators lipman2023flow proposed flow matching (FM), an efficient and simulation-free approach to training CNFs.
 Concretely, they use a per-sample interpolation to derive a simpler objective for learning the marginal vector field that generates a desired probability path in a CNF.
 This formulation provides equivalent gradients without explicit knowledge of the (generally intractable) marginal vector field.
-This work has been extended to different geometries [7, 20] and various applications [52, 9, 12, 21].
-Similar work has been proposed concurrently in [27, 1].
+This work has been extended to different geometries chen2024flow ; klein2023equivariant and various applications wildberger2024flow ; dao2023flow ; ebrahimy2010stock ; kohler2023flow .
+Similar work has been proposed concurrently in liu2022flow ; albergo2023stochastic .
 This paper identifies a reformulation of flow matching that we refer to as variational flow matching (VFM).
@@
 This setting leads to a simple method that we refer to as CatFlow, in which the objective reduces to training a classifier over end points on a per-component basis.
-We apply CatFlow to a set of graph generation tasks, both for abstract graphs [30] and molecular generation [34, 18].
+We apply CatFlow to a set of graph generation tasks, both for abstract graphs martinkus2022spectre and molecular generation ramakrishnan2014quantum ; irwin2012zinc .
 By all metrics, our results match or substantially exceed those obtained by existing methods.
@@
 2.1 Transport Framework for Generative Modeling and CNFs
-Common generative modeling approaches such as normalizing flows [36, 32] and diffusion models [14, 46] parameterize a transformation $\varphi$ from some initial tractable probability density $p_{0}$ – typically a standard Gaussian distribution – to the target data density $p_{1}$.
+Common generative modeling approaches such as normalizing flows rezende2015variational ; papamakarios2021normalizing and diffusion models ho2020denoising ; song2020score parameterize a transformation $\varphi$ from some initial tractable probability density $p_{0}$ – typically a standard Gaussian distribution – to the target data density $p_{1}$.
 In general, there is a trade-off between allowing $\varphi$ to be expressive enough to model the complex transformation while ensuring that the determinant term is still tractable.
@@
 In CNFs, this vector field is learned using a neural network $v_{t}^{\theta}$.
-Through the change of variables formula $p_{t}(x)$ can be evaluated (see appendix D) and hence one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution.
+Through the change of variables formula $p_{t}(x)$ can be evaluated (see appendix˜E) and hence one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution.
 However, obtaining a gradient sample for the loss requires ones to solve the ODE induced during training, making this approach computationally expensive.
 2.2 Flow Matching
-In flow matching [24], the aim is to regress the underlying vector field of a CNF directly on the interval $t\in[0,1]$.
+In flow matching lipman2023flow , the aim is to regress the underlying vector field of a CNF directly on the interval $t\in[0,1]$.
 Flow matching leverages the fact that even though we do not have access to the actual underlying vector field – which we denote as $u_{t}$ – and probability path $p_{t}$, one can construct a per-example formulation by defining conditional flows, i.e.
@@
 First, we can define the marginal vector field and its approximation in terms of an expectation with respect to a distribution over end points of the transformation.
-This implies that we can map a flow matching problem onto a variational counterpart, opening up the usage of variational inference techniques.
+This implies that we can map a flow matching problem onto a variational counterpart.
 Second, under typical assumptions on the forward process, we can decompose the expected conditional vector field into components for individual variables which can be computed in terms of the marginals of the distribution over end points of the conditional trajectories.
@@
 This results in a closed-form objective to train CNFs for categorical data, which we refer to as CatFlow.
-Note that we develop the theory of VFM in section 3 and we relate VFM to flow matching and score-based diffusion in section 4.
+We develop the theory of VFM in section˜3 and we relate VFM to flow matching and score-based diffusion in section˜4.
 3.1 Flow Matching using Variational Inference
-In any flow matching problem, the vector field in eq. 2 can be expressed as an expectation
+In any flow matching problem, the vector field in eq.˜2 can be expressed as an expectation
 $u_{t}(x)$ $=\int u_{t}(x\mid x_{1})p_{t}(x_{1}\mid x)\mathrm{d}x_{1}=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right],$ (6)
@@
 Note that to compute $u_{t}(x)$, one has to evaluate a joint integral over $D$ dimensions.
-This observation leads us to propose a change in perspective.
-Rather than predicting the vector field directly, we can define an approximate vector field in terms of an expectation with respect to a variational distribution $q_{t}^{\theta}$ with parameters $\theta$,
-$v_{t}^{\theta}(x):=\int u_{t}(x\mid x_{1})q_{t}^{\theta}(x_{1}\mid x)\mathrm{d}x_{1}.$ (8)
+This observation leads us to propose a change in parameterization of the learned vector field.
+Rather than predicting the components of the vector field directly, we can define an approximate vector field in terms of an expectation with respect to a variational distribution $q_{t}^{\theta}$ with parameters $\theta$,
+$v_{t}^{\theta}(x):=\int u_{t}(x\mid x_{1})\>q_{t}^{\theta}(x_{1}\mid x)\>\mathrm{d}x_{1}.$ (8)
 Clearly, in this construction $v_{t}^{\theta}(x)$ will be equal to $u_{t}(x)$ when $q_{t}^{\theta}(x_{1}\mid x)$ and $p_{t}(x_{1}\mid x)$ are identical.
 This implies that we can map a flow matching problem onto a variational inference problem.
-Concretely, we can define a variational flow matching problem by minimizing the Kullback-Leibler (KL) divergence from $p_{t}$ to $q_{t}^{\theta}$, which we can express as (see section A.1 for derivations)
+Concretely, we can define a variational flow matching problem by minimizing the Kullback-Leibler (KL) divergence from $p_{t}$ to $q_{t}^{\theta}$, which we can express as
 $\mathbb{E}_{t}\left[\text{KL}\big{(}p_{t}(x)p_{t}(x_{1}\mid x)~||~p_{t}(x)q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right]+\text{const},$ (9)
-where $t\sim\text{Uniform}(0,1)$ and $x,x_{1}\sim p_{t}(x,x_{1})$.
-This leads us to propose the variational flow matching (VFM) objective
+where $t\sim\text{Uniform}(0,1)$ and $x,x_{1}\sim p_{t}(x,x_{1})$ (see section˜A.1 for derivations). This leads us to propose the variational flow matching (VFM) objective
 $\mathcal{L}_{\text{VFM}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right].$ (10)
-Note that, though different, this objective is similar to a typical negative log-likelihood objective.
-At this stage of the exposition, it is not yet clear how useful this variational perspective on flow matching is in practice.
-First, while we can in principle map any flow matching problem onto a variational inference problem, this requires learning an approximation of a potentially complex, high-dimensional distribution $p_{t}(x_{1}\mid x)$.
-Second, we express $v^{\theta}_{t}(x)$ as an expectation, which could well be intractable.
-This could mean that we would have to approximate this expectation, for example using a Monte Carlo estimator.
-We will address these concerns in section 3.2.
+While this variational formulation of flow matching is promising, two potential drawbacks emerge in practical applications.
+First, although it is feasible to reformulate any flow matching problem as a variational inference task, doing so requires learning an approximation of a potentially complex, high-dimensional distribution $p_{t}(x_{1}\mid x)$ (including any correlations between the different $x_{1}^{d}$).
+Second, representing $v^{\theta}_{t}(x)$ as an expectation may pose intractability challenges.
+Interestingly, under typical assumptions about the conditional velocity field in flow matching, this objective simplifies significantly, making it no more computationally demanding than standard flow matching, a point we will further address in section˜3.2.
 3.2 Mean-Field Variational Flow Matching
 Decomposing the conditional vector field.
-At first glance we do not seem to obtain much from this variation view due to the intractability of $p_{t}(x_{1}\mid x)$ and $v_{t}^{\theta}(x)$.
-Fortunately, we can simplify the objective and the calculation of the marginal vector field under certain conditions on the conditional vector field.
-Specifically, we will focus on the special case where the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$, such as in straight line interpolations commonly used in flow matching.
-Informally, we show that in that case the specific form of the distribution used does not affect the expected value of the conditional vector field as long as the marginals coincide.
-As such, the solution can be found by a mean-field/factorized distribution exactly.
-This reduces the problem from one high-dimensional problem, into a series of low dimension problems.
+At first glance, we do not seem to obtain much from this variational view due to the intractability of $p_{t}(x_{1}\mid x)$ and $v_{t}^{\theta}(x)$.
+Fortunately, we can simplify the objective and the calculation of the marginal vector field under the typical case where the conditional vector field $u_{t}(x\mid x_{1})$ is linear in $x_{1}$, such as in straight line interpolations commonly used in flow matching.
+This leads to two simplifications.
+First, we notice that the expected value of a linear conditional velocity field simply equals the conditional velocity field towards the expectation/mean of the distribution we are considering, i.e.
+$u_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]=u_{t}(x\mid\mathbb{E}_{p_{t}}\left[x_{1}\mid x\right])\text{ if $u_{t}(x\mid x_{1})$ is linear in $x_{1}$}.$ (11)
+This means that as long as our variational distribution has the same mean as $p_{t}(x_{1}\mid x)$, we will learn the same underlying field.
+Second, we realize that the expected value of $x_{1}^{d}$ only depends on $p_{t}(x_{1}^{d}\mid x)$, and thus as such as long as $q_{t}^{\theta}(x_{1}\mid x)$ has the same marginal expectations $\mathbb{E}\left[x_{1}^{d}\mid x\right]$, we will learn the same field as under $p_{t}(x_{1}\mid x)$.
+In other words, it suffices to learn a fully-factorized approximation $q_{t}^{\theta}(x_{1}\mid x)$, there is no need to fully characterize the covariance of the posterior probability path.
+Thich allows us to reduce a high-dimensional variational problem to a series of low dimension problems.
 Formally, the following holds:
@@
 $\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right].$ (12)
-We provide a proof in section A.3. It follows directly from theorem 1 that without loss of generality we can consider the considerably easier task of a fully-factorized approximation
+We provide a proof in section˜A.3.
+It follows directly from theorem˜1 that without loss of generality we can consider the considerably easier task of a fully-factorized approximation
 $q_{t}^{\theta}(x_{1}\mid x):=\prod_{d=1}^{D}q_{t}^{\theta}(x_{1}^{d}\mid x).$ (13)
@@
 $\mathcal{L}_{\text{MF-VFM}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right]=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\log q_{t}^{\theta}(x_{1}^{d}\mid x)\right].$ (14)
+Even though we use a factorized model, it is worth emphasizing that due to the linearity of the conditional field, a mean-field variational distribution can learn the solution exactly.
 Computing the marginal vector field.
-Now, to calculate the vector field $v_{t}^{\theta}(x)$, we can simply substitute the factorized distribution $q^{\theta}_{t}(x_{1}\mid x)$ into eq.
-8.
+To calculate the vector field $v_{t}^{\theta}(x)$, we can simply substitute the factorized distribution $q^{\theta}_{t}(x_{1}\mid x)$ into eq.˜8.
 However, this still requires an evaluation of an expectation.
-Fortunately, leveraging the linearity condition significantly simplifies this computation.
-Informally, under this linearity condition, as long as we have access to the first moment of one-dimensional distributions $q^{\theta}_{t}(x_{1}^{d}\mid x)$, we can efficiently calculate $v_{t}^{\theta}(x)$.
-Therefore, if two distributions $q_{t}(x_{1}\mid x)$ and $q_{t}^{\prime}(x_{1}\mid x)$ share the same first moments, they will describe the same vector field, regardless of any other differences they may have.
+Fortunately, leveraging the linearity condition significantly simplifies this computation, since as long as we have access to the first moment of one-dimensional distributions $q^{\theta}_{t}(x_{1}^{d}\mid x)$, we can efficiently calculate $v_{t}^{\theta}(x)$ by simply considering the conditional field towards it.
+Note that for many distributions – e.g.
+Gaussian – learning its parameters is equivalent to learning its expected value.
 Note that the training procedure will differ for two distinct distributions – e.g.
 Gaussian versus Categorical – so the form of the distribution $q^{\theta}_{t}(x_{1}\mid x)$ remains practically important, a flexibility provided through the variational view on flow matching.
-Formally, we can rewrite an arbitrary linear conditional vector field as
-$u_{t}(x\mid x_{1})=A_{t}(x)x_{1}+b_{t}(x),$ (14)
-where $A_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}\times\mathbb{R}^{D}$ and $b_{t}(x):[0,1]\times\mathbb{R}^{D}\rightarrow\mathbb{R}^{D}$.
-If we substitute this into the definition of $v_{t}^{\theta}(x)$ in eq.
-8, we can use the linearity of the expected value to see that
-$v_{t}^{\theta}(x)=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[A_{t}(x)x_{1}+b_{t}(x)\right]=A_{t}(x)\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x).$ (15)
 If we now use the standard flow matching case of using a conditional vector field based on a linear interpolation, the approximate vector field can be expressed in terms of the first moment of the variational approximation:
@@
 Note that this covers both the case of categorical data, which we focus on in this paper, and the case of continuous data, as considered in traditional flow matching methods.
-At first glance, the linearity condition of the conditional vector field $u_{t}(x\mid x_{1})$ in theorem 1 might seem restricting.
+We provided the general algorithm for training and generation in section˜B.1.
+At first glance, the linearity condition of the conditional vector field $u_{t}(x\mid x_{1})$ in theorem˜1 might seem restricting.
 However, in most state-of-the-art generative modeling techniques, this condition is satisfied, e.g.
-diffusion-based models, such as flow matching [24, 2, 27], diffusion models [14, 46], and models that combine the injection of Gaussian noise with blurring [37, 15], among others [10, 43].
-Figure 1: Reparameterized velocities by CatFlow. Note that as $\theta_{t}(x_{t})$ lies in the probability simplex, indeed all velocity implicitly are given the inductive bias of point towards it.
+diffusion-based models, such as flow matching lipman2023flow ; albergo2022building ; liu2022flow , diffusion models ho2020denoising ; song2020score , and models that combine the injection of Gaussian noise with blurring rissanen2022generative ; hoogeboom2022blurring , among others daras2022soft ; singhal2023diffuse .
+Figure 1: Parameterization of the vector field in CatFlow.
+Given an interpolant $x_{t}=tx_{1}+(1-t)x_{0}$, CatFlow predicts a categorical distribution $q^{\theta}_{t}(x_{1}\mid x_{t})$ parameterized by a vector $\mu_{t}(x_{t})$.
+The resulting construction for the vector field $v_{t}^{\theta}(x_{t})=(\mu_{t}(x_{t})-x_{t})/(1-t)$ ensures that trajectories converge to a point on the simplex at $t=1$.
 3.3 CatFlow: Mean-Field Variational Flow Matching for Categorical Data
@@
 In CatFlow, we directly apply the VFM framework to the categorical case.
-Let our parameterised variational distribution $q^{\theta}_{t}(x_{1}^{d}\mid x)=\mathsf{Cat}(x^{d}_{1}\mid\theta_{t}^{d}(x))$, and let us denote $\theta_{t}^{dk}(x):=q_{t}(x_{1}^{d}=k\mid x)$ for brevity.
+Let our parameterised variational distribution $q^{\theta}_{t}(x_{1}^{d}\mid x)=\mathsf{Cat}(x^{d}_{1}\mid\theta_{t}^{d}(x))$, and let us denote the parameters of this categorical as
+$\mu_{t}^{dk}(x):=q^{\theta}_{t}(x_{1}^{d}=k\mid x)=\mathbb{E}_{q^{\theta}_{t}(x_{1}\mid x)}[\mathbb{I}[x_{1}^{d}=k]].$ (16)
 Then, the $d$th component of the learned vector field is
-$v_{t}^{\theta,d}(x):=\sum_{k=1}^{K_{d}}\theta^{dk}_{t}(x)\frac{\mathbb{I}[x_{1}^{d}=k]-x}{1-t}$ (17)
+$v_{t}^{\theta,d}(x):=\sum_{k=1}^{K_{d}}\mu^{dk}_{t}(x)\frac{\mathbb{I}[x_{1}^{d}=k]-x}{1-t}.$ (17)
 Intuitively, CatFlow learns a distribution over the conditional trajectories to all corners of the probability simplices, rather than regressing towards an expected conditional trajectory.
 In the categorical setting, the MF-VFM objective can be written out explicitly. Writing out the probability mass function of the categorical distribution, we see that
-$\log q_{t}^{\theta}(x_{1}^{d}\mid x)=\log\prod_{k=1}^{K^{d}}(\theta_{t}^{dk}(x))^{\mathbb{I}[x_{1}^{d}=k]}=\sum_{d=1}^{D}\mathbb{I}[x_{1}^{d}=k]\log\theta_{t}^{dk}(x).$ (18)
+$\log q_{t}^{\theta}(x_{1}^{d}\mid x)=\log\prod_{k=1}^{K^{d}}(\mu_{t}^{dk}(x))^{\mathbb{I}[x_{1}^{d}=k]}=\sum_{d=1}^{D}\mathbb{I}[x_{1}^{d}=k]\log\mu_{t}^{dk}(x).$ (18)
 As such, we find that CatFlow objective is given by a standard cross-entropy loss:
-$\mathcal{L}_{\text{CatFlow}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\theta^{dk}_{t}(x)\right].$ (19)
+$\mathcal{L}_{\text{CatFlow}}(\theta)=-\mathbb{E}_{t,x,x_{1}}\left[\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\mu^{dk}_{t}(x)\right].$ (19)
 Note, however, that when actually computing $v_{t}^{\theta}$, this can be done efficiently, since
-$\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[u_{t}(x^{d}\mid x_{1}^{d})\right]=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[\frac{x_{1}^{d}-x^{d}}{1-t}\right]=\frac{\theta^{d}(x)-x^{d}}{1-t},$ (20)
-since $\theta^{d}(x):=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}[u_{t}^{d}(x_{1}^{d}\mid x)]$ and the other terms are not in the expectation.
+$\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[u_{t}(x^{d}\mid x_{1}^{d})\right]=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}\left[\frac{x_{1}^{d}-x^{d}}{1-t}\right]=\frac{\mu^{d}(x)-x^{d}}{1-t},$ (20)
+since $\mu^{d}(x):=\mathbb{E}_{q_{t}^{\theta}(x_{1}^{d}\mid x)}[u_{t}^{d}(x_{1}^{d}\mid x)]$ and the other terms are not in the expectation.
 Note that this geometrically corresponds to learning the mapping to a point in the probability simplex, and then flowing towards that.
-This procedure is illustrated in fig.
-1.
+This procedure is illustrated in fig.˜1.
 Because of this, training CatFlow is no less efficient than flow matching.
-More precisely, training CatFlow offers three key benefits over standard flow matching.
-First, the added inductive bias ensures generative paths align with realistic trajectories, improving performance and convergence by avoiding misaligned paths.
-Second, using cross-entropy loss instead of mean-squared error improves gradient behavior during training, enhancing learning dynamics and speeding up convergence.
-Lastly, CatFlow’s ability to learn probability vectors, rather than directly choosing classes, allows it to express uncertainty about variables at a specific time.
+We provided the algorithm for training and generation for CatFlow and a Gaussian VFM objective in section˜B.2 and section˜B.3 respectively.
+More precisely, training CatFlow offers two key benefits over standard flow matching.
+First, given that CatFlow predicts points in the probability simplex and parametrizes the velocities to point into it, an inductive bias is introduced ensuring all generative paths align with realistic trajectories, hence avoiding misaligned paths.
+Second, using a cross-entropy loss instead of a mean-squared error improves gradient behavior during training.
+Both aspects enhance learning dynamics and speed up convergence, which is evaluated in section˜6.3.
+Lastly, CatFlow’s ability to learn probability vectors – rather than directly choosing classes as is common in discrete approaches – allows the model to express uncertainty about variables at a specific time.
 This is especially useful in complex domains like molecular generation, where initial uncertainty about components decreases as more structure is established, leading to more precise predictions.
@@
 Note that even though the following results apply to graphs, an unordered set of categorical variables can be described by a graph without edges.
-Under natural conditions – see appendix A – we ensure CatFlow adheres to this symmetry (see section A.4 for the proof).
+Under natural conditions – see appendix˜A – we ensure CatFlow adheres to this symmetry (see section˜A.4 for the proof).
 Theorem 2.
@@
 Relationship to Flow Matching.
-Informally, the following theorem states that VFM has flow matching as a special case when the variational approximation is Gaussian and under certain assumptions on $u_{t}(x\mid x_{1})$.
-This result is key, as it shows that indeed optimizing a model through VFM is more flexible than regular flow matching.
-Formally, the following holds (see theorem 3 in section A.2 for the proof):
+VFM admits FM as a special case, under certain assumptions on $u_{t}(x\mid x_{1})$, when the variational approximation is Gaussian.
+Formally, the following holds (see theorem˜3 in section˜A.2 for the proof):
 Theorem 3.
@@
 Relationship to Score-Based Models.
-Flow matching [24] is inspired by score-based models [46] and shares strong connections with them.
+Flow matching lipman2023flow is inspired by score-based models song2020score and shares strong connections with them.
+This leads us to two observations.
+The first is that it should be possible to define a variational parameterization of the score function in diffusion models that is analogous to the one in VFM.
+The second is that we can build on existing results that show that many diffusion model objectives, including the standard flow matching objective with a linear interpolant, can be expressed as special cases of a general weighted loss function kingma2021variational ; kingma2023understanding .
+Here we simililarly define a bound on the log-likelihood in terms of a reweighted VFM objective.
 In score-based models, the objective is to approximate the score function $\nabla_{x}\log p_{t}(x)$ with a function $s^{\theta}_{t}(x)$.
-A connection to VFM becomes apparent by observing that the score function can also be expressed as an expectation with respect to $p_{t}(x_{1}\mid x)$ (see section A.5 for derivation):
+A connection to VFM becomes apparent by observing that the score function can also be expressed as an expectation with respect to $p_{t}(x_{1}\mid x)$ (see section˜A.5 for derivation):
 $\nabla_{x}\log p_{t}(x)$ $=\int p_{t}(x_{1}\mid x)\nabla_{x}\log p_{t}(x\mid x_{1})\mathrm{d}x_{1}=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ (22)
@@
 This suggests that there exists a variational formulation of score-based models that is entirely analogous to VFM.
-More specifically, we see that both models could be optimized using a single objective, which yields an approximation $q_{t}^{\theta}(x_{1}\mid x)$ that parameterizes both $v_{t}^{\theta}(x)$ and $s^{\theta}_{t}(x)$.
-Following [46, 3], we can construct stochastic generative dynamics $dx=\tilde{v}_{t}^{\theta}(x)dt+g_{t}dw$ to approximate the true dynamics $dx=\tilde{u}_{t}(x)dt+g_{t}dw$ (see details in section A.6), with
+Indeed, existing work on continuous diffusion for categorical data dieleman2022continuous defines a parameterization of the score function of this form (see section˜5 for a more detailed discussion).
+Following song2020score ; anderson1982reverse , we can construct stochastic generative dynamics $dx=\tilde{v}_{t}^{\theta}(x)dt+g_{t}dw$ to approximate the true dynamics $dx=\tilde{u}_{t}(x)dt+g_{t}dw$ (see details in section˜A.6), with
 $\tilde{u}_{t}(x)$ $:=\mathbb{E}_{p_{t}(x_{1}|x)}\left[u_{t}(x\mid x_{1})+\frac{g_{t}^{2}}{2}\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ $\tilde{v}^{\theta}_{t}(x)$ $:=v_{t}^{\theta}(x)+\frac{g_{t}^{2}}{2}s_{t}^{\theta}(x).$ (24)
@@
 Second, it enables us to show that a reweighted version of the VFM objective provides a bound on the log-likelihood of the model.
-This bound provides another theoretical motivation for learning using the VFM objective.
+This result, inspired by kingma2023understanding , provides another theoretical motivation for learning using the VFM objective.
 Theorem 4.
@@
 where $\lambda_{t}(x)$ is a non-negative function and $C$ is a constant.
-We provide a proof in section A.7.
-We further note that that applying the same linearity condition that we discussed in section 3.2 to the conditional score function maintains all the same connections with score-based models.
+We provide a proof in section˜A.7.
+We further note that that applying the same linearity condition that we discussed in section˜3.2 to the conditional score function maintains all the same connections with score-based models.
 5 Related Work
 Diffusion Models for Discrete Data.
-Several approaches to discrete generation using diffusion models have been developed [4, 28, 53, 17].
-For graph generation specifically, [50] utilize a Markov process that progressively edits graphs by adding or removing edges and altering node or edge categories and is trained using a graph transformer network, that reverses this process to predict the original graph structure from its noisy version.
+Several approaches to diffusion models have been developed for graph generation.
+In vignac2022digress , the authors define a Markov process that progressively edits graphs by adding or removing edges and altering node or edge categories and is trained using a graph transformer network, that reverses this process to predict the original graph structure from its noisy version.
 This approach breaks down the complex task of graph distribution learning into simpler node and edge classification tasks.
-Moreover, [19] proposes a score-based generative model for graph generation using a system of stochastic differential equations (SDEs).
+Moreover, jo2022scorebased proposes a score-based generative model for graph generation using a system of stochastic differential equations (SDEs).
 The model effectively captures the complex dependencies between graph nodes and edges by diffusing both node features and adjacency matrices through continuous-time processes.
-These are non-autoregressive graph generation approaches that perform on par with autoregressive ones, such as in [26, 23, 31].
-Other non-autoregressive approaches worth mentioning are [25, 29, 28].
+These are non-autoregressive graph generation approaches that perform on par with autoregressive ones, such as in liu2018constrained ; liao2019efficient ; mercado2021graph .
+Other non-autoregressive approaches worth mentioning are lippe2020categorical ; madhawa2019graphnvp ; luo2021graphdf .
+There is also work on diffusion-based models for discrete data in the form of text and other sequential data austin2021structured ; luo2021graphdf ; yang2023diffsound ; inoue2023layoutdm .
+Though not explicitly formulated in terms of a variational perspective, the work on continuous diffusion for categorical data dieleman2022continuous arrives at a an approach that is closely related to that of CatFlow.
+This work defines a diffusion in the embedding space of a transformer-based language model.
+It defines an approximation of the score function as an expected value of a conditional score function as in eq.˜23, which leads to an expression for the learned score function in terms of a mean embedding, analogous to the one we obtain in eq.˜15.
+Where this approach differs from CatFlow, other than in that it defines a flow in the embedding space of language models, is in how the objective is defined.
+The authors also minimize a cross-entropy loss, but employ a time-warped objective, similar to the general weighted objective proposed in kingma2021variational , where the warping is optimized to ensure that the entropy of predictions decreases linearly when moving from noise to data in uniform time.
 Flow-based methods for Discrete Data.
-Very recently two flow-based methods for discrete generative modeling have been proposed, which differ both in terms of technical approach and intended use case from the work that we present here.
+Recently, two flow-based methods for discrete generative modeling have been proposed, which differ both in terms of technical approach and intended use case from the work that we present here.
 [^2] Flow matching and diffusion models have also been proposed for geometric graph generation, e.g.
-in [20, 47] and [16, 49] respectively, but since these approaches are continuous (as they generate coordinates based on some conformer) they consider a fundamentally different task than the ones we consider here.
-In [48], a Dirichlet flow framework for DNA sequence design is introduced, utilizing a transport problem defined over the probability simplex.
+in klein2023equivariant ; song2023equivariant and hoogeboom2022equivariant ; trippe2022diffusion respectively, but since these approaches are continuous (as they generate coordinates based on some conformer) they consider a fundamentally different task than the ones we consider here.
+In stark2024dirichlet , a Dirichlet flow framework for DNA sequence design is introduced, utilizing a transport problem defined over the probability simplex, similar to diffusion on simplices proposed in richemond2022categorical .
 This approach differs from CatFlow in that it represents the conditional probability path $p_{t}(x\mid x_{1})$ using a Dirichlet distribution.
 This implies that points $x$ are constrained to the simplex, which is not the case for CatFlow.
-Dirichlet Flows have not been evaluated on graph generation, but we did carry out preliminary experiments based on released source code.
-We have opted not to report these results; we did not obtain good performance out of the box, but have also not invested substantial time in architecture and hyperparameter selection.
-In [6], Discrete Flow Models (DFMs) are introduced.
+Dirichlet Flows have not been evaluated on graph generation, but we did carry out preliminary experiments based on the released source code.
+We compare to this approach in our experiments.
+In campbell2024generative , Discrete Flow Models (DFMs) are introduced.
 DFMs use Continuous-Time Markov Chains to enable flexible and dynamic sampling in multimodal generative modeling of both continuous and discrete data.
@@
 As in the case of Dirichlet flows, no evaluation on graph generation was performed.
-The switch to the variational perspective is inspired by [50], showing significant improvement through viewing the dynamics as a classification task over end points.
+The switch to the variational perspective is inspired by vignac2022digress , showing significant improvement through viewing the dynamics as a classification task over end points.
 However, CatFlow is still a continuous model, and integrates – rather than iteratively samples – during generation.
@@
 Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
-GraphVAE 0.130 0.170 0.050 0.350 0.980 0.540
-GNF 0.030 0.100 0.001 0.200 0.200 0.110
-EDP-GNN 0.052 0.093 0.007 0.053 0.144 0.026
-GDSS 0.021 0.024 0.007 0.045 0.086 0.007
+GraphVAE simonovsky2018graphvae 0.130 0.170 0.050 0.350 0.980 0.540
+GNF liu2019graph 0.030 0.100 0.001 0.200 0.200 0.110
+EDP-GNN niu2020permutation 0.052 0.093 0.007 0.053 0.144 0.026
+GDSS jo2022scorebased 0.021 0.024 0.007 0.045 0.086 0.007
 CatFlow 0.013 0.024 0.008 0.018 0.086 0.007
@@
 We evaluate CatFlow in three sets of experiments.
-First, we consider an abstract graph generation task proposed in [30], where the goal of this task is to evaluate if CatFlow is able to capture the topological properties of graphs.
-Second, we consider two common molecular benchmarks, QM9 [34] and ZINC250k [18], consisting of small and (relatively) large molecules respectively.
+First, we consider an abstract graph generation task proposed in martinkus2022spectre , where the goal of this task is to evaluate if CatFlow is able to capture the topological properties of graphs.
+Second, we consider two common molecular benchmarks, QM9 ramakrishnan2014quantum and ZINC250k irwin2012zinc , consisting of small and (relatively) large molecules respectively.
 This task is chosen to see if CatFlow can learn semantic information in graph generation, such as molecular properties.
 Finally, we perform an ablation comparing CatFlow to standard flow matching, specifically in terms of generalization.
-The experimental setup and model choices are provided in appendix C.
+The experimental setup and model choices are provided in appendix˜D.
 Note that we treat graphs as purely categorical/discrete objects and do not consider ‘geometric’ graphs that are embedded in e.g.
@@
 Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
-MoFlow 91.36 98.65   4.467 63.11 99.99 20.931
-EDP-GNN 47.52 99.25   2.680 82.97 99.79 16.737
-GraphEBM   8.22 97.90   6.143   5.29 98.79 35.471
-GDSS 95.72 98.46   2.900 97.01 99.64 14.656
-Digress 99.00 96.20 - - - -
+MoFlow zang2020moflow 91.36 98.65   4.467 63.11 99.99 20.931
+EDP-GNN niu2020permutation 47.52 99.25   2.680 82.97 99.79 16.737
+GraphEBM liu2021graphebm   8.22 97.90   6.143   5.29 98.79 35.471
+GDSS jo2022scorebased 95.72 98.46   2.900 97.01 99.64 14.656
+Digress vignac2022digress 99.00 96.20 - - - -
+Flow Matching lipman2023flow 94.10 98.20   5.155 94.01 96.68 18.764
+Dirichlet FM stark2024dirichlet 99.10 98.15   0.888 97.52 99.20 14.222
 CatFlow 99.81 99.95   0.441 99.21 100.00 13.211
@@
 We first evaluate CatFlow on an abstract graph generation task, including synthetic and real-world graphs.
-We consider 1) Ego-small (200 graphs), consisting of small ego graphs drawn from a larger Citeseer network dataset [41], 2) Community-small (100 graphs), consisting of randomly generated community graphs, 3) Enzymes (587 graphs), consisting of protein graphs representing tertiary structures of the enzymes from [40], and 4) Grid (100 graphs), consisting of 2D grid graphs.
-We follow the standard experimental setup popularized by [54] and hence report the maximum mean discrepancy (MMD) to compare the distributions of degree, clustering coefficient, and the number of occurrences of orbits with 4 nodes between generated graphs and a test set.
-Following [19], we also use the Gaussian Earth Mover’s Distance kernel to compute the MMDs instead of the total variation.
-The results of the Ego-small and Community-small tasks are summarized in table 1, additional results (and error bars) are provided in appendix B.
+We consider 1) Ego-small (200 graphs), consisting of small ego graphs drawn from a larger Citeseer network dataset sen2008collective , 2) Community-small (100 graphs), consisting of randomly generated community graphs, 3) Enzymes (587 graphs), consisting of protein graphs representing tertiary structures of the enzymes from schomburg2004brenda , and 4) Grid (100 graphs), consisting of 2D grid graphs.
+We follow the standard experimental setup popularized by you2018graphrnn and hence report the maximum mean discrepancy (MMD) to compare the distributions of degree, clustering coefficient, and the number of occurrences of orbits with 4 nodes between generated graphs and a test set.
+Following jo2022scorebased , we also use the Gaussian Earth Mover’s Distance kernel to compute the MMDs instead of the total variation.
+The results of the Ego-small and Community-small tasks are summarized in table˜1, and additional results (and error bars) are provided in appendix˜C.
 The results indicate that CatFlow is able to capture topological properties of graphs, and performs well on abstract graph generation.
@@
 Molecular generation entails designing novel molecules with specific properties, a complex task hindered by the vast chemical space and long-range dependencies in molecular structures.
-We evaluate CatFlow on two popular molecular generation benchmarks: QM9 and ZINC250k [34, 18].
+We evaluate CatFlow on two popular molecular generation benchmarks: QM9 and ZINC250k ramakrishnan2014quantum ; irwin2012zinc .
 We follow the standard setup – e.g.
-as in [42, 28, 50, 19] – of kekulizing the molecules using RDKit [22] and removing the hydrogen atoms.
-We sample 10,000 molecules and evaluate them on validity, uniqueness, and Fréchet ChemNet Distance (FCD) – evaluating the distance between data and generated molecules using the activations of ChemNet [33].
-Here, validity is computed without valency correction or edge resampling, hence following [55] rather than [42, 28], as is more reasonable due to the existence of formal charges in the data itself.
+as in shi2020graphaf ; luo2021graphdf ; vignac2022digress ; jo2022scorebased – of kekulizing the molecules using RDKit Landrum2016RDKit2016_09_4 and removing the hydrogen atoms.
+We sample 10,000 molecules and evaluate them on validity, uniqueness, and Fréchet ChemNet Distance (FCD) – evaluating the distance between data and generated molecules using the activations of ChemNet preuer2018frechet .
+Here, validity is computed without valency correction or edge resampling, hence following zang2020moflow rather than shi2020graphaf ; luo2021graphdf , as is more reasonable due to the existence of formal charges in the data itself.
 We do not report novelty for QM9 and ZINC250k, as QM9 is an exhaustive list of all small molecules under some chemical constraint and all models obtain (close to) 100% novelty on ZINC250k.
 [^3] CatfFlow obtains 49% novelty on QM9.
-The results are summarized in table 2 and samples from the model are shown in fig.
-2.
+The results are summarized in table˜2 and samples from the model are shown in fig.˜2.
 CatFlow obtains state-of-the-art results on both QM9 and ZINC250k, virtually obtaining perfect performance on both datasets.
@@
 (c) Score $5\%$ of data.
-Figure 3: Ablation results, comparison between standard flow matching and CatFlow.
+Figure 3: Ablation results.
+We compare standard flow matching and CatFlow.
+We visualize performance degradation in terms of a score, which is the percentage of molecules that is valid and unique, for a varying number of layers and percentage of the training data.
 To understand the difference in performance between CatFlow and a standard flow matching formulation we perform ablations.
 we focus on generalization capabilities, and as such consider ablations that the number of parameters in the model and the amount of training data.
-In fig.
-3 we report a score, which is the percentage of generated molecules that is valid and unique.
+In fig.˜3 we report a score, which is the percentage of generated molecules that is valid and unique.
 CatFlow not only outperforms regular flow matching in the large model and full data setting, but is also significantly more robust to a decrease in model-size and data.
@@
 In terms of technology readiness, this work is not yet at a level where we foresee direct benefits or risks.
-Acknowledgements
+Acknowledgements.
 This project was supported by the Bosch Center for Artificial Intelligence.
@@
 A.1 Derivation of the Variational Flow Matching Objective
-We derive eq.
-9 that states the equivalence of the optimization of the VFM objective and minimization of the KL divergence between the true endpoint distribution $p_{t}(x_{1}\mid x)$ and the variational approximation $q^{\theta}_{t}(x_{1}\mid x)$.
+We derive eq.˜9 that states the equivalence of the optimization of the VFM objective and minimization of the KL divergence between the true endpoint distribution $p_{t}(x_{1}\mid x)$ and the variational approximation $q^{\theta}_{t}(x_{1}\mid x)$.
 Note that
@@
 $\mathbb{E}_{t,x}\left[\text{H}\big{(}p_{t}(x_{1}\mid x),q^{\theta}_{t}(x_{1}\mid x)\big{)}\right]=-\mathbb{E}_{t,x,x_{1}}\left[\log q_{t}^{\theta}(x_{1}\mid x)\right].$ (30)
-Therefore, the second part of eq. 27 corresponds to negative cross-entropy and minimisation corresponds to maximisation of negative cross-entropy.
+Therefore, the second part of eq.˜27 corresponds to negative cross-entropy and minimisation corresponds to maximisation of negative cross-entropy.
 A.2 Flow Matching as a Special Case of Variational Flow Matching
@@
 $=\mathbb{E}_{t,x,x_{1}}\left[\left\|\left(A_{t}(x)x_{1}+b_{t}(x)\right)-\left(A_{t}(x)\mu_{t}^{\theta}(x)+b_{t}(x)\right)\right\|_{2}^{2}\right]+C$ (34)
-$=\mathbb{E}_{t,x,x_{1}}\left[\left\|u_{t}(x\mid x_{1})-v_{t}^{\theta}(x)\right\|_{2}^{2}\right]+,$ (35)
+$=\mathbb{E}_{t,x,x_{1}}\left[\left\|u_{t}(x\mid x_{1})-v_{t}^{\theta}(x)\right\|_{2}^{2}\right]+C,$ (35)
 which is what we wanted to show. ∎
@@
 $\mathbb{E}_{r}[x]=\mathbb{E}_{p}[x].$ (39)
-Applying this fact to eq. 38 we obtain:
+Applying this fact to eq.˜38 we obtain:
 $\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]$ $=A_{t}(x)\mathbb{E}_{r_{t}(x_{1}\mid x)}\left[x_{1}\right]+b_{t}(x)$ (40)
@@
 Lemma 1.
-If $\theta_{t}(x)$ is permutation equivariant w.r.t $S_{|\mathcal{V}|}$, then so is $v_{t}$.
+If $\mu_{t}(x)$ is permutation equivariant w.r.t $S_{|\mathcal{V}|}$, then so is $v_{t}$.
 Proof.
 Note that
-$v_{t}(\pi\cdot x)=\frac{\theta_{t}(\pi\cdot x)-\pi\cdot x}{1-t}=\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t},$
+$v_{t}(\pi\cdot x)=\frac{\mu_{t}(\pi\cdot x)-\pi\cdot x}{1-t}=\frac{\pi\cdot\mu_{t}(x)-\pi\cdot x}{1-t},$
 where the last step follows from permutation equivariance.
@@
 we conclude that
-$\frac{\pi\cdot\theta_{t}(x)-\pi\cdot x}{1-t}=\pi\cdot\frac{(\theta_{t}(x)-x)}{1-t}=\pi\cdot v_{t}(x),$
+$\frac{\pi\cdot\mu_{t}(x)-\pi\cdot x}{1-t}=\pi\cdot\frac{(\mu_{t}(x)-x)}{1-t}=\pi\cdot v_{t}(x),$
 finishing the proof.∎
 Theorem 5.
-Let $p_{0}$ be an exchangeable distribution – e.g.
-a standard normal distribution – and $\theta_{t}(x)$ be permutation equivariant.
-Then, all permutations of graphs are generated with equal probability.
+Let $p_{0}$ be an exchangeable distribution – e.g. a standard normal distribution – and $\mu_{t}(x)$ be permutation equivariant. Then, all permutations of graphs are generated with equal probability.
 Proof.
-[…]
-[…]
+By result 1, we know that $\mu_{t}(x)$ being permutation equivariant implies that $v_{t}(x)$ is permutation equivariant.
+Moreover, if we let $\Gamma(x):=x+\int_{0}^{1}v_{t}(x_{t})dt$, the for all $\pi\in S_{n}$ we have that
+$\Gamma(\pi\cdot x)=\pi\cdot x+\int_{0}^{1}v_{t}(\pi\cdot x_{t})dt=\pi\cdot x+\int_{0}^{1}\pi\cdot v_{t}(x_{t})dt=\pi\cdot\Gamma(x),$
 where again the last step follows by basic properties of linear operators.
@@
 $\nabla_{x}\log p_{t}(x)$ $=\frac{1}{p_{t}(x)}\nabla_{x}p_{t}(x)$ (43)
-[…]
-[…]
-[…]
-[…]
+$=\frac{1}{p_{t}(x)}\nabla_{x}\int p_{t}(x\mid x_{1})p(x_{1})dx_{1}$ (44)
+$=\frac{1}{p_{t}(x)}\int p(x_{1})\nabla_{x}p_{t}(x\mid x_{1})dx_{1}$ (45)
+$=\int\frac{p_{t}(x\mid x_{1})p(x_{1})}{p_{t}(x)}\nabla_{x}\log p_{t}(x\mid x_{1})dx_{1}$ (46)
+$=\int p_{t}(x_{1}\mid x)\nabla_{x}\log p_{t}(x\mid x_{1})dx_{1}$ (47)
 $=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right].$ (48)
@@
 $u_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[u_{t}(x\mid x_{1})\right]\quad\textrm{and}\quad\nabla_{x}\log p_{t}(x)=\mathbb{E}_{p_{t}(x_{1}\mid x)}\left[\nabla_{x}\log p_{t}(x\mid x_{1})\right],$ (51)
-we can rewrite eq. 50 in the following form:
+we can rewrite eq.˜50 in the following form:
 $dx=\tilde{u}_{t}(x)dt+g_{t}dw,\quad\textrm{where}$ (52)
@@
 Importantly, the function $g_{t}$ does not affect the distribution path $p_{t}(x)$: it only changes the stochasticity of the trajectories.
-In the extreme case when $g_{t}\equiv 0$, the SDE in eq.
-50 coincides with the ODE in eq.
-49.
+In the extreme case when $g_{t}\equiv 0$, the SDE in eq.˜50 coincides with the ODE in eq.˜49.
 Moreover, if we construct the stochastic dynamics in this way, we obtain score-based models as a special case.
@@
 $dx=\left[f_{t}(x)+g_{t}^{2}\nabla_{x}\log p_{t}(x)\right]dt+g_{t}dw.$ (55)
-The SDE in eq. 55 only depends on the score function, so in score-based models, the aim is to learn the score function.
-Now, let us note that similarly to vector field $u_{t}(x)$, the drift term $\tilde{u}_{t}(x)$ in eq.
-52 can be expressed in terms of an end point distribution $p_{t}(x_{1}\mid x)$.
+The SDE in eq.˜55 only depends on the score function, so in score-based models, the aim is to learn the score function.
+Now, let us note that similarly to vector field $u_{t}(x)$, the drift term $\tilde{u}_{t}(x)$ in eq.˜52 can be expressed in terms of an end point distribution $p_{t}(x_{1}\mid x)$.
 Consequently, having a variational approximation of end point distribution $q_{t}^{\theta}(x_{1}\mid x)$, allows us to construct an approximation of the drift term $\tilde{u}_{t}(x)$:
@@
 $dx=\tilde{v}^{\theta}_{t}(x)dt+g_{t}dw\quad\textrm{or}\quad dx=\left[v^{\theta}_{t}(x)+\frac{g_{t}^{2}}{2}s^{\theta}_{t}(x)\right]dt+g_{t}dw.$ (58)
-Then, eq.
-58 is not just simply a new dynamic, it is a family of dynamics that admits deterministic dynamics as a special case when $g_{t}\equiv 0$.
-Importantly, the only thing we need to construct the stochastic process in eq.
-58 is an approximation of the end point distributions $q_{t}^{\theta}(x_{1}\mid x)$.
+Then, eq.˜58 is not just simply a new dynamic, it is a family of dynamics that admits deterministic dynamics as a special case when $g_{t}\equiv 0$.
+Importantly, the only thing we need to construct the stochastic process in eq.˜58 is an approximation of the end point distributions $q_{t}^{\theta}(x_{1}\mid x)$.
 Additionally we know that when $p_{t}(x_{1}\mid x)=q_{t}^{\theta}(x_{1}\mid x)$, $\tilde{u}_{t}(x)=\tilde{v}^{\theta}_{t}(x)$, the processes coincide for all functions $g_{t}$.
@@
 Proof.
-Let us consider the two stochastic processes, that we discussed in section A.6:
+Let us consider the two stochastic processes, that we discussed in section˜A.6:
 $dx=\tilde{u}_{t}(x)dt+g_{t}dw,\quad\textrm{where}\quad dx=\tilde{v}^{\theta}_{t}(x)dt+g_{t}dw.$ (59)
@@
 Next, we can apply Pinsker’s inequality, which states that for two probability distributions $p$ and $q$ the following holds:
-[…]
+$\int|p(x)-q(x)|dx\leq 2\text{KL}(p\|q).$ (66)
 Applying it to the inner integral, we have:
@@
 In this subsection, we discuss the connection between VFM and stochastic dynamics under the condition of linearity in $x_{1}$ of the conditional vector field $u_{t}(x\mid x_{1})$ and conditional score function $\nabla_{x}\log p_{t}(x\mid x_{1})$.
-As we discuss in section 3.2, under the linearity condition, we may express $u_{t}(x)$ in terms of any distribution of end points $r_{t}(x_{1}\mid x)$ if it has the same marginals as $p_{t}(x_{1}\mid x)$.
-In section A.5, we demonstrate that the score function $\nabla_{x}\log p_{t}(x)$ can also be expressed in terms of end point distributions $p_{t}(x_{1}\mid x)$.
+As we discuss in section˜3.2, under the linearity condition, we may express $u_{t}(x)$ in terms of any distribution of end points $r_{t}(x_{1}\mid x)$ if it has the same marginals as $p_{t}(x_{1}\mid x)$.
+In section˜A.5, we demonstrate that the score function $\nabla_{x}\log p_{t}(x)$ can also be expressed in terms of end point distributions $p_{t}(x_{1}\mid x)$.
 Therefore, the score function may also be equally expressed in terms of distribution $r_{t}(x_{1}\mid x)$ if it has the same marginals as $p_{t}(x_{1}\mid x)$.
-This fact is easy to show in the same way as we present in section A.3.
+This fact is easy to show in the same way as we present in section˜A.3.
 Consequently, under the linearity conditions, the drift term $\tilde{u}_{t}(x)$ can also be expressed in terms of $r_{t}(x_{1}\mid x)$, as it is just a linear combination of the vector field $u_{t}(x)$ and score function $\nabla_{x}\log p_{t}(x)$.
-Hence, the transition from the distribution $p_{(}x_{1}\mid x)$ to some distribution $r_{(}x_{1}\mid x)$ does not affect the discussion of connections of stochastic dynamics in section A.6.
-Furthermore, the transition from distribution $p_{t}(x_{1}\mid x)$ to some distribution $r_{t}(x_{1}\mid x)$ does not affect connections of the VFM objective with the model likelihood that we discuss in section A.7.
-It is easy to see that in the derivations, we only rely on functions $\tilde{u}_{t}(x)$ and $\tilde{v}^{\theta}_{t}(x)$ (see eq.
-60).
+Hence, the transition from the distribution $p_{(}x_{1}\mid x)$ to some distribution $r_{(}x_{1}\mid x)$ does not affect the discussion of connections of stochastic dynamics in section˜A.6.
+Furthermore, the transition from distribution $p_{t}(x_{1}\mid x)$ to some distribution $r_{t}(x_{1}\mid x)$ does not affect connections of the VFM objective with the model likelihood that we discuss in section˜A.7.
+It is easy to see that in the derivations, we only rely on functions $\tilde{u}_{t}(x)$ and $\tilde{v}^{\theta}_{t}(x)$ (see eq.˜60).
 However, as we discussed, they are not affected by the transition from $p_{t}(x_{1}\mid x)$ to some $r_{t}(x_{1}\mid x)$.
 Therefore, we may repeat all the same derivations for some $r_{(}x_{1}|x)$ using a factorized distribution.
-Appendix B Detailed Results
+Appendix B Algorithms
+B.1 General Variational Flow Matching
+Algorithm 1 Variational Flow Matching
+# Training
+ Sample $x_{1}$ from data and $x_{0}\sim p_{0}(x_{0})$
+ Sample $t\sim\mathcal{U}(0,1)$
+ Compute $x_{t}=tx_{1}+(1-t)x_{0}$
+ Compute loss $\mathcal{L}=-\log q_{t}^{\theta}(x_{1}\mid x_{t})$
+ Backpropagate.
+# Generation
+ Sample $x_{0}\sim p_{0}(x_{0})$
+ Solve ODE $x_{1}=x_{0}+\int_{t=0}^{t=1}\frac{\mathbb{E}_{q_{t}^{\theta}}\left[x_{1}\mid x_{t}\right]-x_{t}}{1-t+\varepsilon}\text{d}t$
+ Return $x_{1}$
+B.2 Categorical Variational Flow Matching (CatFlow)
+Algorithm 2 Categorical Variational Flow Matching (CatFlow)
+# Assume $q_{t}(x_{1}\mid x_{t})=\prod_{d=1}^{D}\mathsf{Cat}(x^{d}_{1}\mid\mu^{d}_{t}(x_{t}))$ where $\mu$ is learnable
+# Training
+ Sample $x_{1}$ from data and $x_{0}\sim p_{0}(x_{0})$
+ Sample $t\sim\mathcal{U}(0,1)$
+ Compute $x_{t}=tx_{1}+(1-t)x_{0}$
+ Compute cross-entropy loss $\mathcal{L}=-\sum_{d=1}^{D}\sum_{k=1}^{K^{d}}\mathbb{I}[x_{1}^{d}=k]\log\mu^{dk}_{t}(x)$
+ Backpropagate.
+# Generation
+ Sample $x_{0}\sim p_{0}(x_{0})$
+ Solve ODE $x_{1}=x_{0}+\int_{t=0}^{t=1}\frac{\mu_{t}(x_{t})-x_{t}}{1-t+\varepsilon}\text{d}t$
+ Return $x_{1}$
+B.3 Gaussian Variational Flow Matching
+Algorithm 3 Gaussian Variational Flow Matching
+# Assume $q_{t}(x_{1}\mid x_{t})=\prod_{d=1}^{D}\mathcal{N}(x^{d}_{1}\mid\mu^{d}_{t}(x_{t}),I)$ where $\mu$ is learnable
+# Training
+ Sample $x_{1}$ from data and $x_{0}\sim p_{0}(x_{0})$
+ Sample $t\sim\mathcal{U}(0,1)$
+ Compute $x_{t}=tx_{1}+(1-t)x_{0}$
+[…]
+ Backpropagate.
+# Generation
+ Sample $x_{0}\sim p_{0}(x_{0})$
+ Solve ODE $x_{1}=x_{0}+\int_{t=0}^{t=1}\frac{\mu_{t}(x_{t})-x_{t}}{1-t+\varepsilon}\text{d}t$
+ Return $x_{1}$
+Appendix C Additional Results
+C.1 Detailed results
 Here, we provide the results for CatFlow with standard deviations, as computed as in [20] through multiple seeds.
@@
 Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
-$0.013\pm 0.007$ $0.024\pm 0.009$ $0.008\pm 0.005$ $0.018\pm 0.012$ $0.086\pm 0.021$ $0.007\pm 0.005$
+$013\pm 0.007 0.0$$24\pm 0.009 0.0$$08\pm 0.005 0.0$$18\pm 0.012 0.0$$86\pm 0.021 0.0$$007\pm 0.005$
 Enzymes Grid
@@
 Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$ Degree $\downarrow$ Clustering $\downarrow$ Orbit $\downarrow$
-$0.013\pm 0.012$ $0.062\pm 0.011$ $0.008\pm 0.007$ $0.115\pm 0.010$ $0.004\pm 0.002$ $0.075\pm 0.071$
+$013\pm 0.012 0.0$$62\pm 0.011 0.0$$008\pm 0.007$ $115\pm 0.010 0.0$$04\pm 0.002 0.0$$075\pm 0.071$
 QM9 ZINC250k
@@
 Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Unique $\uparrow$ FCD $\downarrow$
-$99.81\pm 0.03$ $99.95\pm 0.02$ $0.441\pm 0.023$ $99.21\pm 0.04$ $100.00\pm 0.00$ $13.211\pm 0.12$
-Appendix C Experimental setup
-C.1 Model
+$9.81\pm 0.03 99$$9.95\pm 0.02$ $0.441\pm 0.023$ $99.21\pm 0.04$ $100.00\pm 0.00$ $13.211\pm 0.12$
+C.2 Extra results SBM and Planar Graphs
+Results on Stochastic Block Model and Planar Graphs.
+We ran extra experiments for (standard) flow matching and Dirichlet flow matching.
+We observe that CatFlow obtains SOTA performance on all tasks and metrics.
+Moreover, it is worth noting CatFlow was significantly faster to train than Dirichlet FM due to a computationally cheaper forward process.
+Model Deg $\downarrow$ Clus $\downarrow$ Orb $\downarrow$ V.U.N. $\uparrow$
+Stochastic Block Model
+SPECTRE [35] 1.9 1.6 1.6 53%
+ConGress [58] 34.1 3.1 4.5 0%
+DiGress [58] 1.6 1.5 1.7 74%
+Flow Matching [27] 10.2 2.0 3.2 22%
+Dirichlet FM [56] 1.7 1.5 1.4 80%
+CatFlow 1.5 1.5 1.4 85%
+Planar Graphs
+SPECTRE [35] 2.5 2.5 2.4 25%
+ConGress [58] 23.8 8.8 2590 0%
+DiGress [58] 1.4 1.2 1.7 75%
+Flow Matching [27] 5.1 5.6 5.5 30%
+Dirchlet FM [56] 1.5 1.3 1.5 80%
+CatFlow 1.4 1.3 1.5 80%
+Appendix D Experimental setup
+D.1 Model
 To ensure a comparison on equal terms to baselines, we employ the same graph transformer network as proposed in [12], which was also used in [58], along with the same hyper-parameter setup.
@@
 $\text{PNA}(X)=\text{cat}(\text{max}(X),\text{min}(X),\text{mean}(X),\text{std}(X))W.$
-C.2 Hyperparameters and Computational Costs
+D.2 Hyperparameters and Computational Costs
 We report the hyperparameters here:
@@
 Learning Rate $2\cdot 10^{-4}$ $2\cdot 10^{-4}$ $2\cdot 10^{-4}$
-Weight Decay $1\cdot 10^{-12}$ $1\cdot 10^{-12}$ $1\cdot 10^{-12}$
+Weight Decay $\cdot 10^{-12} 11$$\cdot 10^{-12}$ $1\cdot 10^{-12}$
 EMA 0.999 0.999 0.999
@@
 All experiments were run on a single NVIDIA RTX 6000 and took about a day to run.
-Appendix D Detail CNFs
+Appendix E Detail CNFs
 To compute the resulting distribution $p_{t}$ for CNF, one can use the change of variables formula:
-$[\varphi_{t}]_{*}p_{0}(x)=p_{0}(\varphi_{t}^{-1}(x))\text{det}\left[\frac{\partial\varphi_{t}^{-1}}{\partial x}(x)\right].$ (71)
-[…]
+$\varphi_{t}]_{*}p_{0}(x)=p_{0}(\varphi_{t}^{-1}(x))\text{det}\left[\frac{\partial\varphi_{t}^{-1}}{\partial x}(x)\right].$ (71)
+This induces a probability path, i.e.
+a mapping $p_{t}:[0,1]\times\mathbb{R}^{D}\to\mathbb{R}_{>0}$ such that $\int p_{t}(x)dx=1$ for all $t\in[0,1]$.
+We say that $v_{t}$ generates this probability path given a starting distribution $p_{0}$.
+In theory, one could try and optimize the empirical divergence between the resulting distribution $p_{1}$ and target distribution, but obtaining a gradient sample for the loss requires us to solve the ODE at each step during training, making this approach computationally prohibitive.
 One way to assess if a vector field generates a specific probability path is using the continuity equation, i.e. we can assess whether $v_{t}$ and $p_{t}$ satisfy
@@
 $x_{1}=x_{0}+\int_{0}^{1}v_{t}(x_{t})\mathrm{d}t.$ (73)
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4. Justification: Our paper follows the structure of the abstract and introduction directly.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: We added a section addressing limitations in section˜7.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory Assumptions and Proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [Yes]
+  14. Justification: We provide all proofs in detail in appendix˜A.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental Result Reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19.
+Justification: Besides code, we provide the model description and hyperparameters in appendix˜D.
+Moreover, our model and datasets are common in the field, e.g.
+[20], so no novel experimental details are required.
+Our main contribution, the objective, is clearly stated in section˜3.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: We provide all code to the experiments. Moreover, all data is publically available, and will be provided in the repository, including data processing.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental Setting/Details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29.
+Justification: We provide the hyperparameters and experimental detail appendix˜D.
+Moreover, we took all of our choices from previous works of [20], which is also publicly available.
+Last, our code will be made public.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment Statistical Significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We provide error bars and how they are computed in appendix˜C.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments Compute Resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: We provide this information in appendix˜D.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code Of Ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We read and checked the code of ethics.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader Impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We reflect on this in our conclusion, i.e. section˜7.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: We do not publish such models and as such do not require safeguards for them.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: We cite the authors of the relevant assets where used, see section˜6. All relevant information about licenses is publicly available and respected in our code.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New Assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [N/A]
+  64. Justification: We provide a new objective to train and use previous code/assets for the experiments which are referenced.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and Research with Human Subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: This work did not involve crowdsourcing nor research with human subjects.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: We did not work with subjects.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.